### v2

# Soft-TransFormers for Continual Learning

Abstract
Inspired by Well-initialized Lottery Ticket Hypothesis (WLTH), which provides suboptimal fine-tuning solutions, we propose a novel fully fine-tuned continual learning (CL) method referred to as Soft-TransFormers (Soft-TF). Soft-TF sequentially learns and selects an optimal soft-network for each task. During sequential training in CL, an well-initialized Soft-TF mask optimizes the weights of sparse layers to obtain task-adaptive soft (real-valued) networks, while keeping the well-pre-trained layer parameters frozen. In inference, the identified task-adaptive network of Soft-TF masks the parameters of the pre-trained network, mapping to an optimal solution for each task and minimizing Catastrophic Forgetting (CF) - the soft-masking preserves the knowledge of the pre-trained network. Extensive experiments on the Vision Transformer (ViT) and the Language Transformer (Bert) demonstrate the effectiveness of Soft-TF, achieving state-of-the-art performance across Vision and Language Class Incremental Learning (CIL) scenarios.
Keywords: Continual Learning (CL), Soft-TransFormers (Soft-TF), Adapter, and LoRA.
1 Introduction
Continual Learning (CL), also known as Lifelong Learning (Thrun, 1995; Rusu et al., 2016; Zenke et al., 2017; Hassabis et al., 2017), is a learning paradigm where a series of tasks are learned sequentially. The principle objective of continual learning is to replicate human cognition, characterized by the ability to learn new concepts incrementally throughout one’s lifespan. An optimal continual learning system could facilitate a positive forward and backward transfer, leveraging the knowledge gained from previous tasks to solve new ones, while also updating its understanding of previous tasks with the new knowledge. However, achieving continual learning is challenging due to the occurrence of catastrophic forgetting or catastrophic interference (McCloskey and Cohen, 1989), a phenomenon where the performance of the model on previous tasks deteriorates significantly when it learns new tasks. This can make it challenging to retain the knowledge acquired from previous tasks, ultimately leading to a decrease in overall performance. To address the issue of catastrophic forgetting during continual learning, numerous conventional approaches have been proposed on Convolutional Neural Networks (CNNs), which can be broadly classified as follows: (1) Regularization-based methods (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020; Titsias et al., 2020; Mirzadeh et al., 2021) aim to keep the learned information of past tasks during continual training aided by sophisticatedly designed regularization terms, (2) Rehearsal-based methods (Rebuffi et al., 2017; Riemer et al., 2018; Chaudhry et al., 2019a, b; Saha et al., 2021) utilize a set of real or synthesized data from the previous tasks and revisit them, and (3) Architecture-based methods (Mallya et al., 2018; Serrà et al., 2018; Li et al., 2019; Wortsman et al., 2020; Kang et al., 2022, 2023, 2024a) propose to minimize the inter-task interference via newly designed architectural components.
Figure 1: Soft-TransFormers (Soft-TF): the objective is to design a fully fine-tuned model that works well across multiple continual learning settings with incurring task-wise soft network training of attention and feedforward networks, leveraged by WLTH.
Developing neural network models that leverage large-scaled pre-trained models. i.e., Vision Transformer (ViT) (Dosovitskiy et al., 2020) and Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) leads to a new paradigm shift referred to as (4) Prompt-based methods in Continual Learning (CL). Prompt-based methods learn continual representations to provide fixed pre-trained transformers with additional instruction. Notably, while L2P (Wang et al., 2022c) stands out as the seminal work that bridges the gap between prompting and continual learning, DualPrompt (Wang et al., 2022b) introduces an innovative approach to affixing complementary prompts to the pre-trained backbone, thereby enabling the acquisition of both task-invariant and task-specific instructions. Additionally, other notable contributions in this field encompass DyTox (Douillard et al., 2022), S-Prompt (Wang et al., 2022a), CODA-P (Smith et al., 2023b), ConStruct-VL (Smith et al., 2023a), ST-Prompt (Pei et al., 2023), and LGCL (Khan et al., 2023). Recently, Qiao et al. (2024) investigated prompt-projection for better generalized continual learners.
With prior developments of representational research, current prompt-based models can be fine-tuned using trainable prompts to improve their performance on sequential tasks, and the fixed pre-trained backbone can consistently provide unforgettable base session knowledge. However, prompt-based models come with several disadvantages and limitations. First, the prompt-based CL model is highly dependent on the quality and design of the sample or the task-relevant prompts. Poorly trained prompts could lead to suboptimal performance or tend to be biased. Second, managing and maintaining a large set of prompts can become cumbersome and unmanageable as the number of tasks increases. Lastly, prompt tuning is not as flexible as full fine-tuning. The only prompt-tuning with the pre-trained model cannot capture all the nuances of uncorrelated sequential tasks even though leveraging the frozen well-initialized model pre-trained on large-scale datasets since the only frozen well-initialized model provides global solutions rather than task-specific solutions. These disadvantages motivate us to make informed decisions about when and how to use prompt-based models and explore alternative methods like full fine-tuning for more robust and flexible prompt-based continual learning performance.
To overcome the limitations of conventional prompt-based methods, the central focus of this work is to pinpoint the parameter-efficient fine-tuning representations of frozen pre-trained networks, comparing with conventional LLM-based fine-tuning methods such as Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2021), as shown in Figure 2. We focus on two main issues when sequential full fine-tuning the pre-trained foundation models: (1) Catastrophic Forgetting (CF) and (2) parameter-efficient fine-tuning CL model. To deploy a practical model to deal with the two points, we suggest a new paradigm for Continual Learning (CL), named Well-initialized Lottery Ticket Hypotehesis:
Well-initialized Lottery Ticket Hypothesis (WLTH). A well-initialized dense neural network contains globally minimal solutions that can retain the prior class knowledge while providing room to learn the new class knowledge through isolated fine-tuning of the networks or subnetworks.
Leveraged by the WLTH, this work proposes a new Soft-TransFormer (Soft-TF) to address fine-tuning with minimal CF, as shown in Figure 1. In this work, we find task-specific soft-networks based on well-trained frozen transformer parameters that incrementally learn task-adaptive weights associated with sequential tasks.
Our contributions can be summarized as follows:
  • Inspired by Well-initialized Lottery Ticket Hypothesis (WLTH), we propose a novel continual learning method referred to as Soft-TransFormers (Soft-TF), which learns compact task-specific soft-networks from well pre-trained parameters for each task.
  • Extensive experiments demonstrate that Soft-TF leads to better generalized parameter-efficient continual models than baselines such as DualPrompts, Adaptor, and LoRA, achieving state-of-the-art performances on Vision and Language Class-Incremental Learning (CIL) scenarios.
Refer to caption Refer to caption Refer to caption
(a) Adapter (b) LoRA (c) Soft-TransFormer
Figure 2: Comparisions of LLM-based fine-tuning methods (Adapters, LoRA) with Soft-TransFormer (Soft-TF): well-initialized Soft-TF $\bm{\theta}\odot\bm{m}_{t}$ is trained at $l_{e}$-th or only few attention layers while LLM fine-tuning methods such as Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2021) are trained at all attention layers. Note that all pre-trained parameters $\bm{\theta}$ are fixed in fine-tuning t-th task parameter $\bm{m}_{t}$ in Continual Learning (CL) scenarios.
2 Related Works
Continual Learning (McCloskey and Cohen, 1989; Thrun, 1995; Kumar and Daume III, 2012; Li and Hoiem, 2016) is the challenge of learning a sequence of tasks continuously while utilizing and preserving previously learned knowledge to improve performance on new tasks. Four major approaches have been proposed to tackle the challenges of continual learning, such as catastrophic forgetting. One such approach is (1) Regularization-based approaches (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020; Titsias et al., 2020; Mirzadeh et al., 2021), which aim to reduce catastrophic forgetting by imposing regularization constraints that inhibit changes to the weights or nodes associated with past tasks. (2) Rehearsal-based approaches (Rebuffi et al., 2017; Chaudhry et al., 2019a, b; Saha et al., 2021; Deng et al., 2021; Sun et al., 2023; Sarfraz et al., 2023; Mai et al., 2021; Lin et al., 2023; Aljundi et al., 2019; Caccia et al., 2021; Chaudhry et al., 2019c; Liang and Li, 2024; Buzzega et al., 2020) store small data summaries to the past tasks and replay them during training to retain the acquired knowledge. Some approaches in this line of work (Shin et al., 2017; Aljundi et al., 2019) accommodate the generative model to construct the pseudo-rehearsals for previous tasks. (3) Architecture-based approaches (Mallya et al., 2018; Serrà et al., 2018; Li et al., 2019; Wortsman et al., 2020; Kang et al., 2022, 2023, 2024b, 2024a) use the additional capacity to expand (Xu and Zhu, 2018), dynamic representation (Yan et al., 2021; Singh et al., 2020) or isolate (Rusu et al., 2016) model parameters, preserving learned knowledge and preventing forgetting. Rehearsal and architecture-based methods have shown remarkable efficacy in suppressing catastrophic forgetting but require additional capacity for the task-adaptive parameters (Wortsman et al., 2020) or the replay buffers. Recently, (4) Prompt-based approaches, an emerging transfer learning technique, harnesses a fixed function of pre-trained Transformer models. This empowers the language model to receive additional instructions for enhancing its performance on downstream tasks. Two key works stand out in the prompt-based CL: L2P (Wang et al., 2022c), a seminal study bridging the gap between prompting and continual learning, and DualPrompt (Wang et al., 2022b), which proposes an innovative method for attaching complementary prompts to a fixed pre-trained backbone. Here, for the first time, we introduce a new approach to update the fixed pre-trained parameters through learnable sparse layered soft-networks under the convergence theory, maximumly enabling the acquisition of task-invariant and task-specific instructions with prompts.
Prompt-based CL. With advances in Vision Transformers (Khan et al., 2022) and prompt-based fine-tuning in NLP (Li and Liang, 2021) and CL (Razdaibiedina et al., 2023), Wang et al. (2022c) have shown that interacting with a pre-trained model via prompt learning is a promising approach, L2P  (Wang et al., 2022c) for continual learning. Then, other advanced prompt-based CL models emerged such as DualPrompt (Wang et al., 2022b), DyTox (Douillard et al., 2022), S-Prompt (Wang et al., 2022a), CODA-P (Smith et al., 2023b), ConStruct-VL (Smith et al., 2023a), ST-Prompt (Pei et al., 2023), and LGCL (Khan et al., 2023). Recently, Prompt Gradient Projection (PGP) (Qiao et al., 2024), a small set of learnable orthogonal parameters, is appended to the input and enables quick adaptation of a frozen ImageNet pre-trained model to new streaming tasks. Their analysis shows that directly leveraging the pre-trained vision-language model without introducing any learnable parameters is a simple yet promising approach to continual learning. The PGP was also adopted to a joint vision-language model like CLIP (Radford et al., 2021) for continual learning, which presents multiple advantages - catering to practical scenarios without well-defined task identities and boundaries in class incremental learning scenarios. However, prompt-based models come with several disadvantages. Poorly trained prompts could lead to suboptimal performance or tend to be biased. Moreover, prompt tuning could not capture all nuances of uncorrelated sequential tasks. These disadvantages lead to exploring alternative methods like full fine-tuning or hybrid approaches for more robust and flexible prompt-based model performance. In this work, to alleviate these issues of prompt tuning, we investigate a fully fine-tuning of well-pre-trained transformers on training soft networks on various continual learning scenarios, comparing with LLM-based CL baselines such as MBPA++ (de Masson D’Autume et al., 2019), IDBR (Huang et al., 2021), and SLM (Bohao et al., 2024).
3 Prerequisites
We start with conventional prompt-based continual learning methods using Vision Transformer (ViT) (Dosovitskiy et al., 2020) in Class Incremental Learning (CIL) scenarios.
3.1 Problem Statement
Continual Learning (CL) involves training deep neural networks (DNN) on time-variant data represented as a sequence of tasks, $\mathcal{D}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{\mathcal{T}}\}$. Each $t$-th task, $\mathcal{D}_{t}=\{(\bm{x}_{i}^{t},y_{i}^{t})^{n_{t}}_{i=1}\}$ consists of $n_{t}$ tuples where $\bm{x}_{i}^{t}\in\mathcal{X}_{t}$ is an input sample and $y_{i}^{t}\in\mathcal{Y}_{t}$ is the corresponding label. When a task $\mathcal{X}_{t}$ arrives, a model $f_{\theta}$ is trained for the current task, while data from previous tasks is inaccessible. This work focuses primarily on class incremental learning (CIL), in which the task-ID is not given during inference.
3.2 Prompt-based Class Incremental Learning (CIL)
A simple yet effective prompt-based (prompt-tuning) CIL model: Learning to Prompt (L2P) (Wang et al., 2022c) is first proposed. In this model, a prompt $p$, a tiny set of trainable tokens combined with image features, is fed into the Vision Transformer (ViT) to help the model resist forgetting. To select suitable prompts for task-specific training, L2P utilizes a prompt pool $P$ containing numerous prompt-key pairs, $\{p_{t},k_{t}\}_{t=1}^{\mathcal{T}}$, where ${p}_{t}\in\mathbb{R}^{1\times D}$ represents the $t$-th task prompt, $k_{t}$ represents the $t$-th coresponding task key, and $\mathcal{T}$ is the total number of prompt-key pairs. Building on L2P, DualPrompt (Wang et al., 2022b) divided the prompts into expert (E-) prompts and general (G-) prompts for distinct features learning. DualPrompt also replaced prompt-tuning with prefix-tuning, which was successfully proven in Language CIL scenarios.
4 Transformer with Learnable Soft-networks
In this section, we explain how Soft-TransFormers (Soft-TF) leverage learnable soft-networks to train sequential tasks while keeping the well-pretrained model parameters fixed. To introduce our novel Soft-TF and provide a clearer understanding, we draw on a partial explanation of DualPrompt.
4.1 Soft-MSA Layers
To address the task-specific fine-tuning of the pre-trained model, such as ViT, this work proposes a new Soft-TransFormer (Soft-TF), as illustrated in Figure 1. The proposed Soft-TF consists of a conventional neural network, like a multilayer transformer with multihead attention and forward networks. Using well-trained transformer parameters, we could discover task-specific soft-networks, as depicted in Figure 4. The Soft-TF incrementally learns model weights and task-adaptive soft-masks with well-pre-trained and soft-network parameters $\bm{m}$.
The Soft-TF originates from learned parameters distributed with $\mu\approx 1.0$ & various variances, as stated in Figure 3: the histogram density plots compare the distribution of parameters for the attention layer L[12] in two contexts: the masking mechanism ($\bm{m}^{QKV}$) and the fine-tuned weighted product ($\bm{w}^{QKV}\odot\bm{m}^{QKV}$) for two tasks (Task 2 and Task 3). In Figure 3(a), the $\bm{m}^{QKV}$ parameters are concentrated in a narrow range centered around 1. There is a slight variation between Task 2 and Task 3. However, the overall range remains consistent. $\bm{m}^{QKV}$ acts as a soft gating mechanism, controlling which pre-trained parameters are retained or modified during fine-tuning. In LABEL:{fig:soft_weight_cifar10}(b), the distribution of pre-trained weight ($\bm{w}^{QKV}$) is centered around 0, with a Gaussian-like spread. The distribution of soft-masked fine-tuned weight ($\bm{w}^{QKV}\odot\bm{m}^{QKV})$ remain similar to that of $\bm{w}^{QKV}$ but slightly narrower. The overlap between the two distributions indicates that the fine-tuning process maintains the general structure of the pre-trained weights while introducing task-specific modifications. By using soft-masks $\bm{m}^{QKV}$, Soft-TF ensures task-specific fine-tuning, reducing the risk of catastrophic forgetting and improving performance in various sequential tasks. The task-specific Soft-TF fine-tuning enables the models to infer task ID ideally in following Vision CIL scenarios.
Refer to caption Refer to caption
(a) $\bm{m}^{QKV}$ (b) $\bm{w}^{QKV}$ v.s. $\bm{w}^{QKV}\odot\bm{m}^{QKV}$
Figure 3: Attention Layer-(L[12]) Histogram Density Estimates of Pre-trained weight ($\bm{\theta}$) and Soft-TF ($\bm{m}$) on 10-Split-CIFAR100.
Given a pre-trained parameter $\bm{\theta}$ and learnable soft-parameters $\bm{m}$, Soft-ViT is represented as $f_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$, consisting of $N$ consecutive soft-MSA layers. We extend the notation by denoting the input embedding feature of the $l_{\ast}$-th learnable soft-MSA layer as $\bm{h}^{(l_{\ast})}$, where $l_{\ast}=1,2,\dots,N$, and $l_{\ast}$ can refer to either the G-Prompt layer $l_{g}$ or the E-Prompt layer $l_{e}$. Note that while the pre-trained parameters $\bm{\theta}$ remain fixed, the soft-parameters $\bm{m}$ are updated to provide task-specific solutions.
G-prompt. $\bm{g}\in\mathbb{R}^{L_{g}\times D}$ with sequence length $L_{g}$ and embedding dimension $D$, is a shared parameter for all tasks. G-Prompt is attached to the $l_{g}$-th MSA layer to transform $\bm{h}^{(l_{g})}$ via a prompting function as follows:
$\bm{h}_{g}^{(l_{g})}=f^{prompt}_{\bm{\theta}}(\bm{g},\bm{h}^{(l_{g})}),$
where $f^{prompt}_{\bm{\theta}}$ defines the approach for attaching the prompt to the hidden embeddings.
E-prompt & Soft-networks. $\bm{e}=\{\bm{e}_{t}\}^{\mathcal{T}}_{t=1}$ is a set of task-dependent parameters, where $\bm{e}_{t}\in\mathbb{R}^{L_{e}\times D}$ has as sequence length of $L_{e}$ and the same embedding dimension $D$ as the G-prompt, and $\mathcal{T}$ is the total number of tasks. Unlike the shared G-prompt, each $\bm{e}_{t}$ is associated with a task-specific key $\bm{k}_{t}\in\mathbb{R}^{D}$, which is also a learnable parameter aimed at capturing representative features of a task. For an input example from the $t$-th task, to attach E-prompt to the $l_{e}$-th soft-MSA layer, we apply the prompting function in a similar way:
$\bm{h}_{e}^{(l_{e})}=f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}(\bm{e}_{t},\bm{h}^{(l_{e})}).$
Figure 4: Soft-TransFormers (Soft-TF): At training time, the E-Prompt and the Soft-network are selected according to task identity, and the selected G-Prompt, E-Prompt, and the Soft-networks (Soft-Attention and Feed Forwards) are trained together with a classifier. At test time, an input is transformed by a query function (Prompt ID) or task identifier (Gradient ID) to match the closest task key $\bm{k}_{t}$, E-prompt $\bm{e}_{t}$ and Soft-networks $\bm{m}_{t}^{\{K,Q,V\}}$. Note task identifier is depicted in Section 5.
4.2 Prompts with Learnable Subnetworks
G- and E-prompts, along with learnable soft-networks, encode specific types of instructions during training with the backbone and work together to guide the model’s predictions during inference. We have demonstrated the method for attaching prompts and learnable soft-networks to a single soft-MSA layer. Similarly to the approach taken in DualPrompt (Wang et al., 2022b), we also investigate layers of E-prompts with learnable soft-networks $\bm{m}$, while utilizing the layers designated for G-prompts.
Layers of G- and E-Prompts. We use the multilayered extension of both types of prompts: $\bm{g}=\{\bm{g}^{(l_{g})}\}_{l_{g}=start_{g}}^{end_{g}}$, where $\bm{g}^{(l_{g})}\in\mathbb{R}^{L_{g}\times D}$ represents the G-prompt attached to the $l_{g}$-th MSA layer. Similarly, we define $\bm{e}_{t}=\{\bm{e}_{t}^{(l_{e})}\}_{l_{e}=start_{e}}^{end_{e}}$ for the $l_{e}$-th conventional MSA layer. In this configuration, the G-prompt $\bm{g}^{(l_{g})}$ is attached from the $start_{g}$-th to the $end_{g}$-th conventional MSA layers, and the E-prompt $\bm{e}_{t}^{(l_{e})}$ is attached to the $\left[start_{e},end_{e}\right]$-th soft-MSA layers, ensuring that there is no overlap between them. In our experiments, we follow the $l_{g}\notin\left[start_{e},end_{e}\right]$ ($l_{g}=[1,2]$) settings used in DualPrompt and empirically search for the optimal $\left[start_{e},end_{e}\right]$ layers for the learnable subnetworks through ablation studies.
Learnable Soft-networks. The prompting function $f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}$ determines how prompts ($\bm{p}$) are combined with fine-tuned soft ($\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}$) embedding features. From another perspective, $f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}$ directly influences the interaction between high-level instructions in the prompts and low-level representations. Therefore, we believe that a well-designed prompting function, along with task-specific parameters, is crucial for optimizing overall continual learning performance.
Specifically, applying a prompting and fine-tuning function $f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}$ can be seen as modifying the inputs to the soft-MSA layers. Let the input to the soft-MSA layer be $\bm{h}\in\mathbb{R}^{L\times D}$, and denote the input query, key, and values for the soft-MSA layer as $\bm{h}_{Q}$, $\bm{h}_{K}$, and $\bm{h}_{V}$, respectively. A soft-MSA layer is defined by the following equation:
$\text{MSA}(\bm{h}_{Q},\bm{h}_{K},\bm{h}_{V})=[\bm{h}_{1},\cdots,\bm{h}_{i},\cdots,\bm{h}_{n}]\bm{w}^{O}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{O}}$
where $\bm{h}_{i}=\text{Attention}(\bm{h}_{Q}(\bm{w}^{Q}_{i}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{Q}}),\bm{h}_{K}(\bm{w}^{K}_{i}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{K}}),\bm{h}_{V}(\bm{w}^{V}_{i}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{V}}))$; $\bm{w}_{i}^{O},\bm{w}_{i}^{Q},\bm{w}_{i}^{K}$, and $\bm{w}_{i}^{V}$ are fixed projection matrices while ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{O}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{Q}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{K}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{V}$} are learnable parameters. $s$ is the number of heads. In ViT, $\bm{h}_{Q}=\bm{h}_{K}=\bm{h}_{V}$. Here, we define a unified prompt parameter with a sequence length of $L_{p}$, such as $\bm{p}\in\mathbb{R}^{L_{p}\times D}$ for a single-layered G- or E-prompt.
4.3 Fine-tuning on Well-initialized Parameters
In this framework, we concatenate the prompts $\bm{p}_{t}$ and the embedding sequence $\bm{x}_{t}$, i.e., inputs from $t$-th task, along the embedding dimension: $\bm{z}_{t}=[\bm{p}_{t};\bm{x}_{t}]$. With the weights of $\bm{w}^{Q}\odot\bm{m}^{Q},\bm{w}^{K}\odot\bm{m}^{K},\bm{w}^{V}\odot\bm{m}^{V}$, the soft-transformer takes query ($\bm{q}_{t}=(\bm{w}^{Q}\odot\bm{m}^{Q})\bm{z}_{t}$) and key ($\bm{k}_{t}=(\bm{w}^{K}\odot\bm{m}^{K})\bm{z}_{t}$) as input of the soft-MSA layer. The soft-attention matrix is then given by:
$\bm{a}_{t}=\text{softmax}\left(\frac{\bm{q}_{t}\bm{k}_{t}^{T}}{\sqrt{D/n}}\right)$
where we focus on $\bm{q}_{t}\bm{k}_{t}^{T}=(\bm{w}^{Q}\odot\bm{m}^{Q})\bm{z}_{t}$ and $\bm{z}_{t}^{T}(\bm{w}^{K}\odot\bm{m}^{K})^{T}$. First, the trainable prompt parameters can be denoted as:
$\bm{z}_{t}\cdot\bm{z}_{t}^{T}=\begin{bmatrix}\bm{p}_{t}\\ \bm{x}_{t}\end{bmatrix}\begin{bmatrix}\bm{p}_{t}\;\bm{x}_{t}\end{bmatrix}=\begin{bmatrix}\bm{p}_{t}\bm{p}^{T}_{t}\;\bm{p}_{t}\bm{x}_{t}^{T}\\ \bm{x}_{t}\bm{p}_{t}^{T}\;\bm{x}_{t}\bm{x}_{t}^{t}\end{bmatrix}$
Second, the trainable soft-attention layer’s parameters with $\bm{m}^{Q}$ and $\bm{m}^{K}$ are as follows:
$\left\{\begin{matrix}\bm{m}^{Q}\cdot\bm{p}_{t}\bm{p}_{t}^{T}\cdot(\bm{m}^{K})^{T},\\ \bm{m}^{Q}\cdot\bm{x}_{t}\bm{p}_{t}^{T}\cdot(\bm{m}^{K})^{T},\\ \bm{m}^{Q}\cdot\bm{p}_{t}\bm{x}_{t}^{T}\cdot(\bm{m}^{K})^{T},\\ \bm{m}^{Q}\cdot\bm{x}_{t}\bm{x}_{t}^{T}\cdot(\bm{m}^{K})^{T}.\\ \end{matrix}\right\}$
where $\bm{w}^{Q}$ and $\bm{w}^{K}$ are frozen and unchanged during training and test. The trainable soft-attention layer’s parameters $\bm{m}^{\ast}$ are used to analyze the convergence rate of the Soft-Transformer (see Section A.1 stated in detail).
4.4 Optimization of Soft-Transformers
The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm 1 and Algorithm 2. We denote the architecture with attached prompts as $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$. The input $\bm{x}$ of the $t$-th task is transformed using $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$ and then passed to the classification head $f\phi$, parameterized by $\phi$, for prediction. Finally, we train the two prompts, the task keys, the soft-attention parameters, and the newly-initialized classification head in an end-to-end manner:
$\min_{\bm{g},\bm{e}_{t},\bm{m}_{t},\bm{k}_{t},\bm{\phi}}\mathcal{L}(f_{\phi}(f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}(\bm{x})),y)+\lambda\mathcal{L}_{\text{match}}(\bm{x},\bm{k}_{t}),$
Here, $\mathcal{L}$ represents the cross-entropy loss, and $\mathcal{L}_{\text{match}}(\bm{x},\bm{k}_{t})=\gamma(\bm{q}(\bm{x}),\bm{k}_{t})$ denotes the matching loss, where $\bm{q}(\bm{x})=f(\bm{x})[0]$ corresponds to the feature vector associated with the [class] token (Dosovitskiy et al., 2020; Wang et al., 2022b), and $\gamma$ is the cosine similarity. The scalar $\lambda$ serves as a balancing factor between the losses; here, we follow the same DualPrompt setting as a baseline.
Table 1: Performances of Vision Class Incremental Learning (CIL, ViTs) in terms of Accuracy, Forgetting, Trainable parameters, and Time complexity on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R. Note that Soft-TF performed without FeedForward ($F$) networks. $\ast$ denotes our reproduced results.
ViT-B/12 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
Method #Tr.Params. Task ID ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec]
L2P 0.03M Prompt ID 83.77 / 6.63 12.00K / 75 71.29 / 13.96 11.50K / 76 60.44 / 9.00 12.80K / 46
 + PGP 0.03M Prompt ID 84.34 / 5.59 12.30K / 75 76.12 / 13.26 12.70K / 76 61.40 / 8.03 13.07K / 46
 + Soft-TF-L[3,4,5] 6.93M Prompt ID 86.26 / 4.79 12.85K / 78 76.17 / 15.77 13.98K /100 69.80 / 5.13 14.23K / 49
 + Soft-TF-L[10,11,12] 6.93M Gradient ID 86.46 / 4.87 12.87K / 128 77.67 / 15.84 13.95K / 160 69.56 / 5.28 14.23K / 79
DualPrompt 0.03M Prompt ID 86.50 / 5.77 12.12K / 76 82.98 / 8.20 11.60K / 78 68.13 / 4.46 13.10K / 47
 + PGP 0.03M Prompt ID 86.92 / 5.35 12.21K / 76 83.74 / 7.91 13.12K / 78 63.34 / 4.53 13.33K / 47
 + Adapter-L[all], r=75 6.93M Prompt ID 86.51 / 4.75 14.98K / 90 84.48 / 5.81 16.13K / 109 70.56 / 4.71 15.65K / 54
 + LoRA-L[10,11,12], r=500 6.93M Prompt ID 82.00 / 4.33 13.24K / 79 92.14 / 2.02 15.89K / 105 43.51 / 13.21 15.09K / 53
 + Soft-TF-L[10,11,12] 6.93M Prompt ID 92.35 / 2.98 13.87K / 80 97.40 / 0.57 15.60K / 104 76.62 / 5.30 15.35K / 52
 + Soft-TF-L[10,11,12] 6.93M Gradient ID 98.05 / 0.25 13.71K / 130 98.96 / 0.23 15.51K / 163 83.70 / 0.53 15.08K / 82
5 Experiments
We validate our Soft-TF on several benchmark datasets against continuous learning baselines in (1) Class-Incremental Learning (CIL) based on ViT models  [^1] https://github.com/ihaeyong/Soft-TF.git and (2) CIL based on language models  [^2] https://github.com/ihaeyong/LLM-Soft-TF.git.
5.1 Experimental Settings
Datasets. (1) CL based on ViT models: we evaluate our method mainly on 1) 10/20-Split-CIFAR100 (Krizhevsky et al., 2009), constructed by splitting the 100 classes into 10 tasks/20 tasks. 2) 10-Split-TinyImageNet (Abai and Rajmalwar, 2019), constructed by splitting the 200 classes into 10 tasks. 3) 10-Split-ImageNet-R (Hendrycks et al., 2021), constructed by splitting the 200 classes into 10 tasks. To show our effectiveness, we additionally compare our method with the baselines on 5-Split-CUB200 and 10-Split-TinyImageNet. (2) CL based on language models: we test our method on the widely adopted continual learning benchmarks for language models following de Masson D’Autume et al. (2019), which use five text classification datasets (Chen et al., 2020; Zhang et al., 2015) including AG News (news classification), Yelp (sentiment analysis), DBPedia (Wikipedia article classification), Amazon (sentiment analysis) and Yahoo Answers (Q&A classification).
Baselines. To validate the powerfulness of our method in (1) CL based on ViT models, we compare our results with various CIL baselines including, L2P (Wang et al., 2022c), DualPrompt (Wang et al., 2022b), and DualPrompt-PGP (Qiao et al., 2024). Moreover, to show the effectiveness of our method in (2) CL based on language models, we set MBPA++ (de Masson D’Autume et al., 2019), IDBR (Huang et al., 2021), SLM (Bohao et al., 2024), and DualPrompt (Wang et al., 2022b) as baselines.
Table 2: Performances of Language Class Incremental Learning (CIL, Bert) in terms of Accuracy, Forgetting, Trainable parameters, and Time complexity on Bert benchmark datasets. “DR”: data replay. $\ast$ denotes our reproduced results.
Order: 4 Order: 5 Order: 6 Order: 7 Average Bert-base (110.0M)
Method DR Task ID ACC/Forget ACC/Forget ACC/Forget ACC/Forget ACC/Forget #\#Tr.Params Tr./Test[sec]
MBPA++ $\checkmark$ L.Adapter 74.9 / 3.71 73.1 / 2.76 74.9 / 2.89 74.1 / 3.00 74.3 / 3.09 110.0M 30.12K / 87.43
IDBR $\checkmark$ L.Adapter 75.9 / 3.46 76.2 / 2.33 76.4 / 2.88 76.7 / 2.90 76.3 / 2.89 110.0M 27.73K / 86.22
SLM Retrieval 78.2 / 0.70 77.8 / 0.70 78.0 / 0.70 78.2 / 0.70 78.1 / 0.70 7.32M 17.51K / 73.36
DualPrompt-L[3,4,5] Prompt ID 69.7 / 0.02 67.5 / 0.04 69.1 / 0.04 68.8 / 0.03 68.9 / 0.03 0.24M 13.86K / 70.14
DualPrompt-L[10,11,12] Prompt ID 67.2 / 0.06 65.1 / 0.09 66.5 / 0.08 66.3 / 0.07 66.3 / 0.08 0.24M 13.86K / 70.14
+ Adapter-L[all], r=16 Prompt ID 75.0 / 0.00 75.1 / 0.00 75.1 / 0.00 75.1 / 0.00 75.1 / 0.00 7.54M 19.45K / 75.34
+ LoRA-L[all], r=96 Prompt ID 78.9 / 0.00 78.9 / 0.00 78.9 / 0.00 78.9 / 0.00 78.9 / 0.00 7.28M 21.57K / 76.61
+ Soft-TF-L[3,4,5] (SOTA) Prompt ID 79.3 / 0.00 79.6 / 0.00 79.4 / 0.00 79.4 / 0.00 79.4 / 0.00 7.23M 15.51K / 71.12
+ Soft-TF-L[10,11,12] Prompt ID 79.5 / 0.00 79.3 / 0.00 79.3 / 0.00 79.5 / 0.00 79.4 / 0.00 7.23M 15.53K / 71.12
Task Inference. At the inference time, we infer task identity for arbitrary pieces of task samples $\bm{x}$ for finding the proper task nuances and demonstrating full fine-tuning results. We summarize the following two methods:
  • Prompt ID: For a test example $\bm{x}$, we simply choose the best matched task index via $\text{argmin}_{t}\gamma(q(\bm{x}),\bm{k}_{t})$.
[…]
Table 3: Performances of Vision Class Incremental Learning (CIL, ViTs) in terms of Pretained-dataset (ImageNet-21K, SAM, DINO) and Task-IDs on 10-Split-CIFAR100 and 5-Split-CUB200.
10-Split-CIFAR100 5-Split-CUB200
Method Pretrained-dataset Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
DualPrompt ImageNet-21K Prompt ID 86.50 5.77 82.02 4.23
DualPrompt-PGP ImageNet-21K Prompt ID 86.92 5.35 82.46 3.76
DualPrompt SAM Prompt ID 86.11 6.08 82.02 4.73
DualPrompt DINO Prompt ID 64.18 23.81 50.88 10.10
 +Soft-TF ImageNet-21K Prompt ID 92.42 2.44 76.17 9.04
 +Soft-TF (SOTA) ImageNet-21K Gradient ID 97.87 0.21 87.93 0.66
 +Soft-TF (SOTA) SAM Gradient ID 97.87 0.21 87.93 0.66
 +Soft-TF DINO Gradient ID 84.50 12.27 69.79 10.93
Upper-Bound of Soft-TF - - 93.90 - 85.56 -
Refer to caption Refer to caption
(a) L2P-PGP v.s. Soft-Transformer (b) DualPrompt v.s. Soft-Transformer
Figure 5: Layer-wise(L[$\ast$]) Performances of Soft-TF on 10-Split-CIFAR100. Note that L[9,10,11] denotes Soft-TransFormer of 9, 10, 11 Layers.
5.2 Performances
Performances of Soft-TF on Vision CIL. We compare our Soft-TransFormers (Soft-TF) with state-of-the-art CIL baselines, as shown in Table 1. Our Soft-TF significantly outperforms all baselines: Adapter and LoRA and upper-bounds of Soft-TF, including L2P and DualPrompt, in accuracy and forgetting measurements. The time complexity of training and testing does not seem to be a significant issue under the same trainable parameters. The performance gain of Soft-TF is especially notable in DualPrompt-based learning compared to L2P, suggesting the importance of global prompt-tuning and multi-head attention prompt-tuning in DualPrompt. Additionally, task-identity inference using Gradient-ID is crucial for achieving full fine-tuning results in CIL.
Performances of Soft-TF on Language CIL. Our Soft-TF significantly outperforms all baselines including MBPA++, IDBR, SLM, and DualPrompt in accuracy and forgetting measurements, as shown in Table 2. Replay-based methods, i.e., MBPA++ and IDBR, carry a computational burden since they replay the previous task’s samples and fine-tune 110.0M parameters. SLM shows inefficient time complexity since it takes two stages (warmup and fine-tuning) in the training step and the retrieval process in the inference step. Moreover, Soft-TF performs better than LLM-based fine-tuning baselines such as Adapter and LoRA under the same trainable parameter settings. Soft-TF trains and tests faster than LLM-based baselines since Soft-TF fine-tunes sparse layers (i.e., [3,4,5] layers). Note that LoRA failed to fine-tune onto sparse [3,4,5] layers.
5.3 Ablation Studies
Well-initialized LTH (WLTH) on Vision CIL. To demonstrate the efficacy of our proposed method on Well-initialized Lottery Ticket Hypothesis (WLTH) backbones, we evaluate our Soft-TransFormers (Soft-TF) by extending two distinct pre-trained models, ViT-DINO and ViT-SAM (Caron et al., 2021; Chen et al., 2021). As shown in Table 3, we tested our method on the 10-Split-CIFAR100 and 5-Split-CUB200 datasets using three pre-trained ViTs: ImageNet-21K, DINO, and SAM, further validating the effectiveness of our approach on non-ImageNet datasets (Krizhevsky et al., 2009; Wah et al., 2011). Surprisingly, when initialized with ImageNet-21K and SAM, DualPrompt-Soft-TF with ImageNet-21K achieved the same performance levels. Moreover, DualPrompt-Soft-TF outperformed all baselines, i.e., DualPrompt-PGP, on both benchmark datasets, indicating that well-initialized weights provide better generalization in continual learning scenarios.
Random initialization. Random initialization of Soft-Transformer’s weights plays a critical role when leveraging well-pretrained models like Vision Transformers (ViTs). The optimal training point is the parameters of a well-pretrained model. Among the initialization methods, Uniform initialization for Soft-TransFormer satisfies this requirement effectively. To validate these claims, we analyze the impact of common random initialization methods, including Xavier, Kaiming, Normal, and Uniform Initialization, as shown in Table 4. The results demonstrate that the same well-initialization point leads to independent optimal task performance, particularly with Gradient ID inference. Furthermore, this ablation study strengthens our Soft-TF with state-of-the-art-perfomances inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH).
Table 4: Random initialized Performances of Class Incremental Learning (CIL, ViTs) in terms of accuracy and forgetting on 10-Split-CIFAR100. Note ”w/o FF” denotes ”Soft fine-tuning without FeedForward (FF)” networks. DualPrompt’s parameters are pretrained on ImageNet-21K.
10-Split-CIFAR100
DualPrompt Task ID Random Init. ACC($\uparrow$) Forget($\downarrow$)
 +Soft-TF-L[10,11,12] w/o FF Prompt ID Xavier 90.59 3.85
 +Soft-TF-L[10,11,12] w/o FF Prompt ID Kaiming 90.72 3.63
 +Soft-TF-L[10,11,12] w/o FF Prompt ID Normal 90.45 3.78
 +Soft-TF-L[10,11,12] w/o FF Prompt ID Uniform(1.0, 1.0) 92.35 2.98
 +Soft-TF-L[10,11,12] w/o FF Gradient ID Uniform(1.0, 1.0) 98.05 0.25
Upper-Bound of Soft-TF 93.90 -
Layer-wise Inspections. We analyze the layer-wise performance of Soft-Transformer with respect to L2P and DualPrompt on the 10-Split-CIFAR100 dataset to identify the optimal configurations, as shown in Figure 5. Our observations reveal that the global prompt in DualPrompt influences Soft-Transformer’s performance differently in L2P and DualPrompt settings. In L2P-PGP, the best performance was achieved with Soft-TransFormers applied to the lower layers ((a) L2P-Soft-TF-L[1,2]-PGP), whereas in DualPrompt, the higher layers ((b) DualPrompt-Soft-TF-L[10,11,12]) yielded the best results. Notably, DualPrompt-Soft-TF-L[10,11,12] without PGP demonstrated impressive performance, achieving almost zero forgetting (0.21). These findings suggest that our approach could significantly enhance the effectiveness of large-scale Transformer models in continual learning scenarios.
Refer to caption Refer to caption
(a) DualPrompt-L[12] (b) Soft-TF-L[12]
Figure 6: Attention map comparisons of DualPrompt and Soft-TF: at Layer-(L[12]) on a single category 7 sample $\bm{x}_{i}^{t}$ of 10-Split-CIFAR100.
Attention Layer. We compare attention maps from the 12th attention layer (L[12]) of two different models, DualPrompt and Soft-TF, applied to category 7 of the 10-Split-CIFAR100 dataset, as shown in Figure 6. While DualPrompt distributes its attention more uniformly, Soft-TF prioritizes specific patches of the input image, likely those more relevant to Category 7. These attention maps highlight the differences in how DualPrompt and Soft-TF allocate attention in a Vision Transformer. Soft-TF’s sharper focus on essential features aligns with the reported superior performance on the 10-Split-CIFAR100 dataset. This visualization emphasizes the interpretability and task-specific optimization of Soft-TF over the more generalized behavior of DualPrompt.
6 Conclusion
Inspired by Well-initialized Lottery Ticket Hypothesis (WLTH), which provides suboptimal fine-tuning solutions, we proposed a novel fully fine-tuned continual learning (CL) method referred to as Soft-TransFormers (Soft-TF). Soft-TF sequentially learns and selects an optimal soft-network for each task. During sequential training in CL, an well-initialized Soft-TF mask optimizes the weights of sparse layers to obtain task-adaptive soft (real-valued) networks, while keeping the well-pre-trained layer parameters frozen. In inference, the identified task-adaptive network of Soft-TF masks the parameters of the pre-trained network, mapping to an optimal solution for each task and minimizing Catastrophic Forgetting (CF) - the soft-masking preserves the knowledge of the pre-trained network. Extensive experiments on the Vision Transformer (ViT) and the Language Transformer (Bert) demonstrated the effectiveness of Soft-TF, achieving state-of-the-art performance across Vision and Language CIL scenarios.
Appendix A Appendix
A.1 Analysis of Soft-TransFormers (Soft-TF)
Analysis of Soft-TransFormers for Convex-Lipschitz Functions. To analyze the convergence rate of the Soft-TransFormers (Soft-TF), we limit ourselves to the case of convex-Lipshitz functions along with the analysis (Shalev-Shwartz and Ben-David, 2014). Let $\bm{w}^{\ast}=\{\bm{g}^{\ast},\bm{e}_{t}^{\ast},\bm{m}_{t}^{\ast}\}$ be any vector or an optimal solution and let $B$ be an upper bound on $\bm{w}^{\ast}||$ when $\bm{w}^{(1)}=\bm{0}$. It is convenient to think of $\bm{w}^{\ast}$ as the minimizer of $f(\bm{w})$, but the analysis that follows holds for every $\bm{w}^{\ast}$.
We obtain an upper bound on the sub-optimality of our solution with respect to $\bm{w}^{\ast}$, namely, $f(\bar{\bm{w}})-f(\bm{w}^{\ast})$, where $\bar{\bm{w}}=\frac{1}{T}\bm{w}^{(t)}$. From the definition of $\bar{\bm{w}}$, and using Jensen’s inequality, we have that
$\begin{aligned} f(\bar{\bm{w}})-f(\bm{w}^{\ast})&=f\left(\frac{1}{T}\sum_{t=1}^{T}\bm{w}^{(t)}\right)-f(\bm{w}^{\ast})\\ &\leq\frac{1}{T}\sum_{t=1}^{T}\left(f(\bm{w}^{(t)})\right)-f(\bm{w}^{\ast})\\ &=\frac{1}{T}\sum_{t=1}^{T}\left(f(\bm{w}^{(t)})-f(\bm{w}^{\ast})\right).\end{aligned}$
For every $t$, because of the convexity of $f$, we have that
$f(\bm{w}^{(t)})-f(\bm{w}^{\ast})\leq\left<\bm{w}^{(t)}-\bm{w}^{\ast},\nabla f(\bm{w}^{(t)})\right$>
Combining the preceeding we obtain
$f(\bm{w}^{(t)})-f(\bm{w}^{\ast})\leq\frac{1}{T}\sum_{t=1}^{T}\left<\bm{w}^{(t)}-\bm{w}^{\ast},\nabla f(\bm{w}^{(t)})\right$>
To bound the right-hand side we rely on the following lemma:
Lemma 1.
Let $\bm{v}_{1},\cdots,\bm{v}_{T}$ be an arbitrary sequence of vectors. Any algorithm with an well initialization (pre-trained model) $\bm{w}^{(1)}\neq\bm{0}$ and an update rule of the form
$\bm{w}^{(t+1)}=\bm{w}^{(t)}-\eta\bm{v}_{t}$
[…]
where $\bm{w}_{m}\neq\bm{w}_{q}$ since $\bm{m}$ is learnable parameters in Soft-TransFormers. Specifically, we could assume that $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$ and $\bm{w}_{p}=(\bm{w}^{Q}\odot\bm{1}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{1}^{K})^{T}$ of Equation 6.
Theorem 2.
[…]
$\frac{1}{T}\sum_{t=1}^{T}\left<\bm{w}^{(t)}-\bm{w}^{\ast},\bm{v}_{t}\right>\leq\frac{B_{m}\rho}{\sqrt{T}}<\frac{B_{p}\rho}{\sqrt{T}}<\frac{B\rho}{\sqrt{T}}.$
Proof  Using algebraic manipulations (completing the square), we obtain:
[…]
where the last equality follows from the definition of the update rule. Summing the equality over $t$, we have
[…]
The first sum on the right-hand side is a telescopic sum that collapses to
[…]
Plugging this in Equation, we have
[…]
where the last equality is due to the definition $\bm{w}^{(1)}=\bm{0}$. This proves the first part of the lemma. The second part follows by upper bounding $\bm{w}||$ by $B$, $\bm{v}_{t}||$ by $\rho$, deciding by $T$, and plugging in the value of $\eta$.
In terms of Soft-TransFormers $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$ of Equation 6, we have
[…]
where $\bm{v}_{t}$ is an arbitrary $t$-th vector and $\bm{w}^{(1)}\neq\bm{0}$ since $\bm{w}^{Q}$ and $\bm{q}^{K}$ are pre-trained parameters.
however, in term of prompt $\bm{w}_{p}=(\bm{w}^{Q}\odot\bm{1}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{1}^{K})^{T}$, we have
[…]
where $\bm{v}_{t}$ is an arbitrary $t$-th vector of prompt and $\bm{w}^{(1)}\neq\bm{0}$ since $(\bm{w}\odot\bm{1})^{Q}$ and $(\bm{w}\odot\bm{1})^{K}$ are pre-trained parameters, $\bm{w}^{Q}$ and $\bm{w}^{K}$, respectively.
[…]
$\frac{1}{T}\sum_{t=1}^{T}\left<\bm{w}^{(t)}-\bm{w}^{\ast},\bm{v}_{t}\right>\leq\frac{B_{m}\rho}{\sqrt{T}}<\frac{B_{p}\rho}{\sqrt{T}}<\frac{B\rho}{\sqrt{T}}.$
Refer to caption Refer to caption
(a) DualPrompt-PGP v.s. Soft-TF (b) DualPrompt-WSN v.s. Soft-TF
[…]
A.2 Experimental Details
For fair comparisons with the baselines (Wang et al., 2022c, b; Qiao et al., 2024), we use ViT B/16 (Dosovitskiy et al., 2020) pre-trained on ImageNet-21K as our image encoder, which is kept frozen during training. We train and test on a single Quadro RTX 8000-48GB GPU for baselines and our Soft-TransFormers with Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
We adhere to the experimental settings outlined by Qiao et al. (2024) to validate our method’s effectiveness. When comparing our approach with L2P-PGP and Soft-Transformer on the 10/20-Split-CIFAR100 and 10-Split-TinyImageNet datasets, we train the network for 5 epochs with a batch size of 16 and set the prompt length to 5. For the 10-Split-ImageNet-R dataset, we use 50 epochs, a batch size of 16, and a prompt length of 30. In comparison with DualPrompt-PGP and Soft-TransFormers on the 10/20-Split-CIFAR100 dataset, we train the network for 20 epochs with a batch size of 24 and set the expert prompt length to 5. For the 10-Split-TinyImageNet dataset, we use 5 epochs, a batch size of 24, and an expert prompt length of 5. For the 10-Split-ImageNet-R dataset, we set the epochs to 50, the batch size to 24, and the expert prompt length to 20. Additionally, in all benchmark data sets, the general prompt length is set to 5, and the location inserted into the prompt is kept consistent.
CLIP on CIL and TIL. We conduct our experiments on the 10-Split-CIFAR100 dataset under both Class Incremental Learning (CIL) and Task Incremental Learning (TIL) settings, as shown in Table 5. The results demonstrate that CLIP-Prompt-Soft-TF-L[1-12] significantly improves performance in both settings, indicating that our Soft-TF with Gradient ID is also effective in vision-language models, thereby broadening its applicability.
Table 5: Comparisons of Soft-TF with baselines based on CLIP model on 10-Split-CIFAR100. $\ast$ denotes our reproduced results.
Class Incremental Task Incremental
Method Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
CLIP Prompt ID 73.76 5.60 92.69 2.34
CLIP-PGP Prompt ID 79.47 4.23 93.00 1.58
CLIP Prompt ID 74.60 7.75 93.59 2.80
CLIP-PGP Prompt ID 74.63 7.76 93.67 2.83
CLIP-Prompt Prompt ID 70.27 12.95 93.36 3.07
CLIP-Prompt-Soft-TF-L[3,4,5] Prompt ID 71.58 7.73 95.29 1.12
CLIP-Prompt-Soft-TF-L[3,4,5] Gradient ID 76.77 5.59 95.29 1.12
CLIP-Prompt-Soft-TF-L[1-12] Prompt ID 72.28 3.44 96.83 0.44
CLIP-Prompt-Soft-TF-L[1-12] (SOTA) Gradient ID 85.90 3.07 96.83 0.44
For CLIP-PGP and Soft-TransFormers, we configure a single trainable image prompt that is shared across all tasks within the vision encoder. For the text encoder, following the approach of Qiao et al. (2024), we set a trainable text prompt for each class, which is only trained on the corresponding task. In our comparisons with CLIP-PGP and Soft-TransFormers on the 10-Split-CIFAR100 dataset, we set the image prompt length to 5, the number of epochs to 5, and the batch size to 32.
Table 6: Performances of Class Incremental Learning (CIL) in terms of accuracy and forgetting on $10/20$-Split-CIFAR100 and 10-Split-ImageNet-R. Exemplar means the total buffer size for rehearsal methods.
10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
Method Exemplar Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
BiC 5,000 - 81.42 17.31 73.02 6.23 64.63 22.25
DER++ 5,000 - 83.94 14.55 - - 66.73 20.67
iCaRL 5,000 - 66.00 5.33 78.02 5.80 - -
DER+MCG 2,000 - 67.62 14.64 65.84 13.72 - -
BiC 1,000 - 66.11 35.24 63.12 21.89 52.14 36.70
DER++ 1,000 - 61.06 39.87 - - 55.47 34.64
iCaRL 1,000 - 61.25 14.19 71.32 15.98 - -
FT - - 33.61 86.87 33.52 53.69 28.87 63.80
EWC - - 47.01 33.27 36.73 35.19 35.00 56.16
LWF - - 60.69 27.77 39.12 57.91 38.54 52.37
L2P - Prompt ID 83.77 6.63 71.29 13.96 60.44 9.00
L2P-PGP - Prompt ID 84.34 5.59 76.12 13.26 61.40 8.03
L2P-PGP-Soft-TF - Prompt ID 86.26 4.79 76.17 15.77 69.80 5.13
L2P-PGP-Soft-TF - Gradient ID 86.46 4.87 77.67 15.84 69.56 5.28
DualPrompt - Prompt ID 86.50 5.77 82.98 8.20 68.13 4.46
DualPrompt-Soft-TF-L[3,4,5] - Prompt ID 91.77 3.37 94.43 2.02 74.70 6.46
DualPrompt-Soft-TF-L[3,4,5] - Gradient ID 93.76 1.83 95.38 1.73 82.15 2.20
DualPrompt-WSN-L[10,11,12], c=80.0 $\%$ - Gradient ID 97.41 0.18 90.25 9.08 74.83 0.91
DualPrompt-WSN-L[10,11,12], c=81.0 $\%$ - Gradient ID 97.50 0.21 96.72 2.21 74.21 1.41
DualPrompt-WSN-L[10,11,12], c=82.0 $\%$ - Gradient ID 97.67 0.27 96.44 1.62 75.02 0.92
DualPrompt-WSN-L[10,11,12], c=87.0 $\%$ - Gradient ID 97.51 0.27 97.68 0.75 77.96 1.02
DualPrompt-WSN-L[10,11,12], c=90.0 $\%$ - Gradient ID 97.46 0.38 98.09 0.65 78.80 0.47
DualPrompt-Soft-TF-L[10,11,12] - Gradient ID 97.87 0.21 99.05 0.24 82.38 0.59
DualPrompt-PGP - Prompt ID 86.92 5.35 83.74 7.91 69.34 4.53
DualPrompt-PGP-Soft-TF-L[3,4,5] - Prompt ID 92.41 2.44 95.14 1.90 74.65 4.39
DualPrompt-PGP-Soft-TF-L[3,4,5] - Gradient ID 92.92 2.34 95.89 1.64 81.45 2.89
Upper-Bound of DualPrompt - - 90.85 - 90.85 - 79.13 -
Upper-Bound of Soft-TF - - 93.90 - 93.90 - 80.21 -
Table 7: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, training, and test time on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R. Note ”w/o FF” denotes ”Soft finetuning without FeedForward (FF)” networks.
Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
DualPrompt # Train Params. Task ID Train (sec.) Test (sec.) Train (sec.) Test (sec.) Train (sec.) Test (sec.)
DualPrompt 0.03M Prompt ID 12.12K 76 11.60K 78 13.10K 47
  PGP 0.03M Prompt ID 12.21K 76 13.12K 78 13.33K 47
  Soft-TF-L[12] w/ only ATTN 1.76M Gradient ID 12.18K 129 13.30K 113 13.35K 65
  Soft-TF-L[12] w/ only ATTN 1.76M Prompt ID 12.18K 78 13.30K 80 13.35K 48
  Soft-TF-L[12] w/o FF 2.31M Gradient ID 12.24K 103 13.40K 132 13.42K 66
  Soft-TF-L[11,12] w/o FF 4.62M Gradient ID 12.95K 115 14.38K 146 14.23K 73
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID 13.71K 130 15.51K 163 15.08K 82
  Soft-TF-L[10,11,12] w/o FF 6.93M Prompt ID 13.87K 80 15.60K 104 15.35K 52
  LoRA-L[10,11,12] w/o FF, r=4 0.06M Prompt ID 11.95K 77 11.71K 79 14.34K 48
  LoRA-L[10,11,12] w/o FF, r=24 0.32M Prompt ID 12.03K 78 15.10K 100 15.02K 50
  LoRA-L[10,11,12] w/o FF, r=500 6.93M Prompt ID 13.24K 79 15.89K 105 15.09K 53
  Adapter-L[10,11,12] w/ FF, r=1 0.09M Prompt ID 12.44K 84 12.40K 81 14.40K 50
  Adapter-L[10,11,12] w/ FF, r=4 0.36M Prompt ID 12.80K 85 15.35K 105 14.68K 51
  Adapter-L[10,11,12] w/ FF, r=300 6.93M Prompt ID 13.66K 88 15.72K 106 15.50K 53
Table 8: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, Accuracy, and Forget on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R. Note ”w/o FF” denotes ”Soft finetuning without FeedForward (FF)” networks.
Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
DualPrompt # Train Params. Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
DualPrompt 0.03M Prompt ID 86.50 5.77 82.98 8.20 68.13 4.46
  PGP 0.03M Prompt ID 86.92 5.35 83.74 7.91 69.34 4.53
  Soft-TF-L[12] w/ only ATTN 1.76M Gradient ID 97.17 0.40 98.09 0.54 72.31 3.94
  Soft-TF-L[12] w/ only ATTN 1.76M Prompt ID 94.59 1.12 96.96 1.02 71.13 4.93
  Soft-TF-L[12] w/o FF 2.31M Gradient ID 96.84 0.55 97.81 0.57 81.18 1.31
  Soft-TF-L[11,12] w/o FF 4.62M Gradient ID 97.58 0.34 98.65 0.43 83.09 0.42
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID 98.05 0.25 98.96 0.23 83.70 0.53
  Soft-TF-L[10,11,12] w/o FF 6.93M Prompt ID 92.35 2.98 97.40 0.57 76.62 5.30
  LoRA-L[10,11,12] w/o FF, r=4 0.06M Prompt ID 82.19 4.33 93.74 2.07 70.91   9.11
  LoRA-L[10,11,12] w/o FF, r=24 0.32M Prompt ID 86.77 4.27 95.65 1.04 69.81 10.30
  LoRA-L[10,11,12] w/o FF, r=500 6.93M Prompt ID 82.00 4.33 92.14 2.02 43.51 13.21
  Adapter-L[10,11,12] w/ FF, r=1 0.09M Prompt ID 86.38 4.87 85.61 5.04 70.95 4.31
  Adapter-L[10,11,12] w/ FF, r=4 0.36M Prompt ID 86.53 4.52 85.66 5.00 70.82 4.90
  Adapter-L[10,11,12] w/ FF, r=300 6.93M Prompt ID 86.45 4.61 84.75 5.11 70.55 4.74
Training & Test Time. To clearly illustrate the time complexity of Soft-TF, we present the training and testing times for 10/20-Split-CIFAR100 and 10-Split-ImageNet-R, as shown in Table 7. As the number of trainable parameters in Soft-TF increases, training and testing time complexities grow accordingly. While the testing time complexity of Gradient ID increased by approximately 1.6 times across the three benchmark datasets, it consistently improved task performance on all benchmarks. The corresponding performance metrics are detailed in Table 8. We also investigate the most parameter-efficient and gradient-based task inference methods, as shown in Table 9 and Table 10. Our findings reveal that the 3-shot Gradient ID inference cost (using samples within a mini-batch) with the last layer (Soft-TF-L[12]) is approximately 1.1 times that of Prompt ID, maintaining comparable efficiency while delivering superior performance. Note that m-batch denotes mini-batch.
Table 9: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, training, and test time on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R. Note ”w/o FF” denotes ”Soft finetuning without FeedForward (FF)” networks.
Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
DualPrompt # Train Params. Task ID Train (sec.) Test (sec.) Train (sec.) Test (sec.) Train (sec.) Test (sec.)
DualPrompt 0.03M Prompt ID 12.12K 76 11.60K 78 13.10K 47
  PGP 0.03M Prompt ID 12.21K 76 13.12K 78 13.33K 47
  Soft-TF-L[12] w/o FF 2.31M Prompt ID 12.24K 79 13.40K 80 13.42K 48
  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 3-shot 12.24K 88 13.40K 90 13.42K 57
  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 5-shot 12.24K 94 13.40K 98 13.42K 61
  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 7-shot 12.24K 95 13.40K 108 13.42K 62
  Soft-TF-L[12] w/o FF 2.31M Gradient ID, m-batch 12.24K 103 13.40K 132 13.42K 66
  Soft-TF-L[10,11,12] w/o FF 6.93M Prompt ID 13.87K 80 15.60K 104 15.35K 52
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 3-shot 13.71K 96 15.51K 106 15.08K 63
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 5-shot 13.71K 96 15.51K 109 15.08K 74
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 7-shot 13.71K 106 15.51K 119 15.08K 75
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, batch 13.71K 130 15.51K 163 15.08K 82
Table 10: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, Accuracy, and Forget on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R. Note ”w/o FF” denotes ”Soft finetuning without FeedForward (FF)” networks.
Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
DualPrompt # Train Params. Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
DualPrompt 0.03M Prompt ID 86.50 5.77 82.98 8.20 68.13 4.46
  PGP 0.03M Prompt ID 86.92 5.35 83.74 7.91 69.34 4.53
  Soft-TF-L[12] w/o FF 2.31M Prompt ID 91.83 2.99 96.43 1.00 72.45 5.32
  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 3-shot 93.12 1.82 96.43 1.00 73.55 4.80
  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 5-shot 96.13 0.58 96.43 1.00 75.04 4.49
  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 7-shot 96.51 0.65 96.43 1.00 76.34 4.75
  Soft-TF-L[12] w/o FF 2.31M Gradient ID, batch 96.84 0.55 97.81 0.57 81.18 1.31
  Soft-TF-L[10,11,12] w/o FF 6.93M Prompt ID 92.35 2.98 97.40 0.57 74.62 5.30
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 3-shot 93.92 1.62 97.40 0.57 74.99 4.21
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 5-shot 97.37 0.53 97.40 0.57 77.40 2.91
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 7-shot 97.76 0.51 97.40 0.57 79.33 3.57
  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, m-batch 98.05 0.25 98.96 0.23 83.70 0.53
Soft & Subnetworks have been explored in continual learning through the following notable approaches such as supermasks (Wortsman et al., 2020), WSN (Kang et al., 2022), SoftNet (Kang et al., 2023), and FSO (Kang et al., 2024a). All prior works have primarily demonstrated the effectiveness of sparse solutions in fully connected networks and Convolutional Neural Networks (CNNs). Sparse fine-tuning mechanisms have not been investigated for pre-trained transformers, such as ViT, and remain unexplored. To discover the competitive sparseness in Transformers, we detail the WSN-style task-adaptive fine-tuning and the learnable soft-networks $\bm{m}$ of Transformers, presenting these adaptations for the first time with empirical observations. Here, the Soft-TF originates from learned parameters distributed with $\mu\approx 1.0$ & various variances, as stated in Figure 3: we observed that each task solution has its unique variances which help infer task-ID in gradient step. To demonstrate the effectiveness of Soft-TF, Figure 7(a) presents a radar chart comparing DualPrompt-PGP and Soft-TransFormers across four benchmark datasets with various sparseness (Figure 7(b) & Table 6.): we compare the performance of DualPrompt-WSN and Soft-TF using accuracy across tasks on the 10-Split CIFAR100 dataset. The performances of Soft-TF provide the upper bound of WSN, as shown in Figure 8(a). In Vision Transformers, Figure 8(b) demonstrates the exsistances of a smooth solution of SoftNet.
Comparisions of Soft-TF with LLMs. To demonstrate the effectiveness of Soft-TF, we compare Soft-TF against LLM fine-tuning methods such as Adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021), as shown in Table 8. Under identical experimental conditions—including trainable model parameters (6.9M per task), layers (L[10,11,12]), and Prompt ID—Soft-TF outperformed other LLM-based fine-tuning approaches. The results highlight that directly updating well-pretrained model parameters and prompt-tuning via Soft-TF is more effective than combining representations through LoRA or learning representations with Adapters. Furthermore, we observed that Soft-TF and the other methods exhibited comparable training and testing time complexity for the same number of trainable parameters. Notably, single-layer fine-tuning using Soft-TF (with L[12]) surpasses the performance of the baselines. These findings firmly establish Soft-TF as the most competitive approach among strong LLM baselines (Adapters and LoRA) in the continual learning (CIL) scenario.
Sparsity of Transformer. We inspect the sparse solution through WSN as shown in Table 6, Table 11, and Table 12. We found a suboptimal sparse solution (c=87.0 % on 10-Split-TinyImageNet) with minimal CF through the inspections, as stated in Table 11. This demonstrates the Lottery Ticket Hypothesis (LTH) in transformers, a competitive sparse subnetwork in DenseNetwork. In addition, DualPrompt is the lower-bound while DualPrompt-Soft-TF-$\ast$ is the upper-bound, close to the optimal performances.
Refer to caption Refer to caption
(a) WSN-TransFormers v.s. Soft-TransFormers (b) WSN v.s. SoftNet
Figure 8: Comparisions of Soft-TransFormers with Subnetworks on 10-Split-CIFAR100. Note that L[10,11,12] denotes the fine-tuning layers of 10, 11, and 12.
Table 11: Performances of Subnetworks (WSN) in Class Incremental Learning (CIL) on 10-Split-TinyImageNet.
TinyImageNet
Method Pretrained-Dataset Task ID ACC($\uparrow$) Forget($\downarrow$)
DualPrompt - Prompt ID 86.50 5.77
DualPrompt-WSN-L[10,11,12] C=80.0% Gradient ID 89.95 0.98
DualPrompt-WSN-L[10,11,12] C=81.0% Gradient ID 89.99 1.06
DualPrompt-WSN-L[10,11,12] C=82.0% Gradient ID 89.59 1.18
DualPrompt-WSN-L[10,11,12] C=83.0% Gradient ID 90.60 0.72
DualPrompt-WSN-L[10,11,12] C=85.0% Gradient ID 90.09 1.07
DualPrompt-WSN-L[10,11,12] C=87.0% Gradient ID 91.91 0.38
DualPrompt-WSN-L[10,11,12] C=90.0% Gradient ID 91.28 0.42
DualPrompt-WSN-L[10,11,12] C=93.0% Gradient ID 91.41 0.40
DualPrompt-WSN-L[10,11,12] C=95.0% Gradient ID 90.91 0.83
DualPrompt-Soft-TF-L[10,11,12] - Gradient ID 97.87 0.21
Table 12: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, Accuracy, and Forget on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R. Note ”w/ only ATTN” denotes ”Soft finetuning only with attention (ATTN)” and FeedForward ($F$) networks.
Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
DualPrompt #Tr. Params. Task ID ACC Forget ACC Forget ACC Forget
DualPrompt 0.03M Prompt ID 86.50 5.77 82.98 8.20 68.13 4.46
  PGP 0.03M Prompt ID 86.92 5.35 83.74 7.91 69.34 4.53
  Soft-TF-L[12] w/ only ATTN 1.76M Gradient ID 97.17 0.40 98.09 0.54 72.31 3.94
  Soft-TF-L[12] w/ only ATTN, c=90% 1.58M Gradient ID 96.81 0.61 97.51 1.68 71.67 3.94
Peudo Codes. The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm 1 and Algorithm 2. We denote the architecture with attached prompts as $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$. The input $\bm{x}$ from the $t$-th task is transformed using $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$ and then passed to the classification head $f_{\phi}$, parameterized by $\phi$, for prediction. Finally, we train the two prompts, the task keys, the soft-attention parameters, and the newly-initialized classification head in an end-to-end manner.
Algorithm 1 DualPrompt-Soft-TF at training time
1:  Input: Pre-trained transformer-based backbone $f$, final classification layer $f_{\phi}$,
2:       number of tasks $\mathcal{T}$, training set $\{\{\bm{x}_{i,t},y_{i,t}\}_{i=1}^{n_{t}}\}_{t=1}^{\mathcal{T}}$, G-Prompt $\bm{g}$, E-Prompt $\bm{E}=\{\bm{e}_{t}\}_{t=1}^{\mathcal{T}}$,
3:       task keys $\bm{K}=\{\bm{k}_{t}\}_{t=1}^{T}$, soft-networks $\bm{M}=\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t}}\}_{t=1}^{\mathcal{T}}$ $start_{g},end_{g},start_{e},end_{e}$,
4:       prompting function $f^{prompt}_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$,
5:       number of training epochs of the $t$-th task $\mathcal{K}_{t}$.
6:  Initialize: $\phi,\bm{g},\bm{E},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{M}},\bm{K}$
7:  for task $t=1,\cdots,\mathcal{T}$ do
8:  Select the task-specific E-Prompt, soft-network $\bm{e}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t}}$ and corresponding task key $\bm{k}_{t}$
9:  Generate the prompted architecture $f_{\bm{g},\bm{e}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t}}}$: attach $\bm{g}$ and $\bm{e}_{t}$ to $start_{g}$-th to $end_{g}$-th
10:       and $start_{e}$-th to $end_{e}$-th soft MSA layers respectively, with $f^{prompt}_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$.
11:  for batch $\bm{e}_{s}\sim\mathcal{K}_{t}$ do
12:   Draw a mini-batch $B=\{(\bm{x}_{i,t},y_{i,t})\}_{i=1}^{l}$
13:   for ($\bm{x},y$) in B do
14:    Calculate the prompted feature by
15:    Calculate the per sample loss $\mathcal{L}_{x}$ via
16:   end for
17:   Update $\phi,\bm{g},\bm{E},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{M}},\bm{K}$ by back-propagation
18:  end for
19:  end for
Algorithm 2 DualPrompt-Soft-TF at test time
1:  Given components: Pre-trained transformer-based backbone $f$, trained
2:       $\bm{K}=\{\bm{k}_{t}\}_{t=1}^{\mathcal{T}},\bm{M}=\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t}}\}_{t=1}^{\mathcal{T}},start_{g},end_{g},start_{e},end_{e},$ prompting function $f^{prompt}_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$
3:  Input: test example $\bm{x}$ from mini-batch $\bm{b}$
4:  Select task inference method: (1) Prompt ID or (2) Gradient ID
5:     (1) Prompt ID:
6:       Generate query feature $q(\bm{x})$
7:       Matching for the index of E-Prompt via $t_{\bm{x}}=\mathrm{argmin}_{t}\gamma(q(\bm{x}),\bm{k}_{t})$
8:     (2) Gradient ID:
9:       Assigning each learned subnetwork $\bm{m}_{t}$ a weight $\alpha_{t}$ such that $\sum_{t}\alpha_{t}=1$ and $\alpha_{t}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1/\mathcal{T}}>0.$
10:       Given $\bm{x}\in\bm{b}$ to classify, we can compute our loss $\mathcal{L}=\mathcal{H}(f^{prompt}_{\bm{\theta}\odot(\sum_{t}\alpha_{t}\bm{m}_{t})}(\bm{x}))$
11:       Matching for the index of E-Prompt via $t_{\bm{x}}=\mathrm{argmin}_{t}\frac{\partial\mathcal{H}}{\partial\alpha_{t}}$
12:  Select the task-specific E-prompt $\bm{e}_{t_{\bm{x}}}$ and learned subnetwork $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t_{\bm{x}}}$}
13:  Generate the prompted architecture $f_{\bm{g},\bm{e}_{t_{\bm{x}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t_{\bm{x}}}}}$:
14:       Attaching $\bm{g}$ and $\bm{e}_{t_{\bm{x}}}$ to $start_{g}$-th to $end_{g}$-th
15:       and $start_{e}$-th to $end_{e}$-th MSA layers respectively, with $f^{prompt}_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$.
16:  Prediction: $f_{\bm{g},\bm{e}_{t_{\bm{x}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t_{\bm{x}}}}}(\bm{x})$
Density of Parameters. We inspect the histogram density estimate of the last (12) layer’s parameters of DualPrompt-Soft-TF: attention of QKV ((a) ATTN.QKV) and Projection ((b) ATTN.PROJ) and multi-layer perception (MLP) of FC1 and FC2, as shown in Figure 9. ATTN’s QKV parameters have the largest variance among the parameter densities, while MLP-FC2’s are the smallest. From this observation, we conclude that fine-tuning ATTN’s QKV is required to achieve optimal task performance. In other words, QKV’s parameters are more critical than others.
Refer to caption Refer to caption
(a) ATTN.QKV ($\bm{m}^{QKV}$) (b) ATTN.PROJ ($\bm{m}^{O}$)
Refer to caption Refer to caption
(c) MLP.FC1 ($\bm{m}^{FC1}$) (d) MLP.FC2 ($\bm{m}^{FC2}$)
Figure 9: Layer-(L[12]) Histogram Density Estimates of DualPrompt-Soft-TF’s Parameters on 10-Split-CIFAR100.
Pre-trained Parameters vs. Soft-TF. We inspect the histogram density estimate of the last (12) layer’s parameters of pre-trained model and DualPrompt-Soft-TF: attention of QKV ((a) ATTN.QKV) and Projection ((b) ATTN.PROJ) and multi-layer perception (MLP) of FC1 and FC2, as shown in Figure 10. The most parameters of DualPrompt-Soft-TF are trained around zero-values. Particularly, the difference between pre-trained model’s parameters and Soft-TF is distinctive at QKV module.
Refer to caption Refer to caption
(a) ATTN.QKV ($\bm{w}^{QKV}$ v.s. $\bm{w}^{QKV}\odot\bm{m}^{QKV}$) (b) ATTN.PROJ ($\bm{w}^{O}$ v.s. $\bm{w}^{O}\odot\bm{m}^{O}$)
Refer to caption Refer to caption
(c) MLP.FC1 ($\bm{w}^{FC1}$ v.s. $\bm{w}^{FC1}\odot\bm{m}^{FC1}$) (d) MLP.FC2 ($\bm{w}^{FC2}$ v.s $\bm{w}^{FC2}\odot\bm{m}^{FC2}$)
Figure 10: Layer-(L[12]) Histogram Density Estimates of Pre-trained Weight and DualPrompt-Soft-TF’s Parameters on 10-Split-CIFAR100.

### v3

@@
-Soft-TransFormers for Continual Learning
+Soft-TransFormer for Continual Learning
 Abstract
-Inspired by Well-initialized Lottery Ticket Hypothesis (WLTH), which provides suboptimal fine-tuning solutions, we propose a novel fully fine-tuned continual learning (CL) method referred to as Soft-TransFormers (Soft-TF).
-Soft-TF sequentially learns and selects an optimal soft-network for each task.
-During sequential training in CL, an well-initialized Soft-TF mask optimizes the weights of sparse layers to obtain task-adaptive soft (real-valued) networks, while keeping the well-pre-trained layer parameters frozen.
-In inference, the identified task-adaptive network of Soft-TF masks the parameters of the pre-trained network, mapping to an optimal solution for each task and minimizing Catastrophic Forgetting (CF) - the soft-masking preserves the knowledge of the pre-trained network.
-Extensive experiments on the Vision Transformer (ViT) and the Language Transformer (Bert) demonstrate the effectiveness of Soft-TF, achieving state-of-the-art performance across Vision and Language Class Incremental Learning (CIL) scenarios.
-Keywords: Continual Learning (CL), Soft-TransFormers (Soft-TF), Adapter, and LoRA.
+Inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH), we introduce Soft-Transformer (Soft-TF), a parameter-efficient framework for continual learning that leverages soft, real-valued subnetworks over a frozen pre-trained Transformer.
+Instead of relying on manually designed prompts or adapters, Soft-TF learns task-specific multiplicative masks applied to the key, query, value, and output projections in self-attention.
+These masks enable smooth and stable task adaptation while preserving shared representations.
+Combined with a lightweight dual-prompt mechanism, Soft-TF maintains strong knowledge retention and mitigates Catastrophic Forgetting (CF).
+Across multiple continual learning benchmarks, Soft-TF achieves state-of-the-art performance, consistently outperforming prompt-based, adapter-based, and LoRA-style baselines while requiring minimal additional parameters.
+Machine Learning, Continual Learning, Sparse TransFormers
 1 Introduction
-Continual Learning (CL), also known as Lifelong Learning (Thrun, 1995; Rusu et al., 2016; Zenke et al., 2017; Hassabis et al., 2017), is a learning paradigm where a series of tasks are learned sequentially.
-The principle objective of continual learning is to replicate human cognition, characterized by the ability to learn new concepts incrementally throughout one’s lifespan.
-An optimal continual learning system could facilitate a positive forward and backward transfer, leveraging the knowledge gained from previous tasks to solve new ones, while also updating its understanding of previous tasks with the new knowledge.
-However, achieving continual learning is challenging due to the occurrence of catastrophic forgetting or catastrophic interference (McCloskey and Cohen, 1989), a phenomenon where the performance of the model on previous tasks deteriorates significantly when it learns new tasks.
-This can make it challenging to retain the knowledge acquired from previous tasks, ultimately leading to a decrease in overall performance.
-To address the issue of catastrophic forgetting during continual learning, numerous conventional approaches have been proposed on Convolutional Neural Networks (CNNs), which can be broadly classified as follows: (1) Regularization-based methods (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020; Titsias et al., 2020; Mirzadeh et al., 2021) aim to keep the learned information of past tasks during continual training aided by sophisticatedly designed regularization terms, (2) Rehearsal-based methods (Rebuffi et al., 2017; Riemer et al., 2018; Chaudhry et al., 2019a, b; Saha et al., 2021) utilize a set of real or synthesized data from the previous tasks and revisit them, and (3) Architecture-based methods (Mallya et al., 2018; Serrà et al., 2018; Li et al., 2019; Wortsman et al., 2020; Kang et al., 2022, 2023, 2024a) propose to minimize the inter-task interference via newly designed architectural components.
+Continual Learning (CL), also known as Lifelong Learning (Thrun, 1994; Rusu et al., 2016; Zenke et al., 2017; Hassabis et al., 2017), aims to endow models with the ability to acquire new knowledge over time without erasing previously learned information.
+An ideal CL system would enable both forward transfer, where prior knowledge facilitates learning new tasks, and backward transfer, where new experiences refine earlier representations.
+However, modern neural networks remain highly susceptible to catastrophic forgetting (McCloskey and Cohen, 1989), where updates from new tasks interfere with parameters essential for earlier ones, resulting in significant performance degradation.
+A large body of prior work - predominantly developed for convolutional neural networks - tackles catastrophic forgetting through three major strategies: (1) Regularization-based methods (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020; Titsias et al., 2020; Mirzadeh et al., 2021) constrain updates to preserve knowledge of previously learned tasks; (2) Rehearsal-based methods (Rebuffi et al., 2017; Riemer et al., 2018; Chaudhry et al., 2019a, b; Saha et al., 2021) store or synthesize past data for replay; and (3) Architecture-based methods (Mallya et al., 2018; Serrà et al., 2018; Li et al., 2019; Wortsman et al., 2020; Kang et al., 2022) reduce inter-task interference via task-specific modules or structural expansions.
 Figure 1: Soft-TransFormers (Soft-TF): the objective is to design a fully fine-tuned model that works well across multiple continual learning settings with incurring task-wise soft network training of attention and feedforward networks, leveraged by WLTH.
-Developing neural network models that leverage large-scaled pre-trained models.
-i.e., Vision Transformer (ViT) (Dosovitskiy et al., 2020) and Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) leads to a new paradigm shift referred to as (4) Prompt-based methods in Continual Learning (CL).
-Prompt-based methods learn continual representations to provide fixed pre-trained transformers with additional instruction.
-Notably, while L2P (Wang et al., 2022c) stands out as the seminal work that bridges the gap between prompting and continual learning, DualPrompt (Wang et al., 2022b) introduces an innovative approach to affixing complementary prompts to the pre-trained backbone, thereby enabling the acquisition of both task-invariant and task-specific instructions.
-Additionally, other notable contributions in this field encompass DyTox (Douillard et al., 2022), S-Prompt (Wang et al., 2022a), CODA-P (Smith et al., 2023b), ConStruct-VL (Smith et al., 2023a), ST-Prompt (Pei et al., 2023), and LGCL (Khan et al., 2023).
-Recently, Qiao et al.
-(2024) investigated prompt-projection for better generalized continual learners.
-With prior developments of representational research, current prompt-based models can be fine-tuned using trainable prompts to improve their performance on sequential tasks, and the fixed pre-trained backbone can consistently provide unforgettable base session knowledge.
-However, prompt-based models come with several disadvantages and limitations.
-First, the prompt-based CL model is highly dependent on the quality and design of the sample or the task-relevant prompts.
-Poorly trained prompts could lead to suboptimal performance or tend to be biased.
-Second, managing and maintaining a large set of prompts can become cumbersome and unmanageable as the number of tasks increases.
-Lastly, prompt tuning is not as flexible as full fine-tuning.
-The only prompt-tuning with the pre-trained model cannot capture all the nuances of uncorrelated sequential tasks even though leveraging the frozen well-initialized model pre-trained on large-scale datasets since the only frozen well-initialized model provides global solutions rather than task-specific solutions.
-These disadvantages motivate us to make informed decisions about when and how to use prompt-based models and explore alternative methods like full fine-tuning for more robust and flexible prompt-based continual learning performance.
-To overcome the limitations of conventional prompt-based methods, the central focus of this work is to pinpoint the parameter-efficient fine-tuning representations of frozen pre-trained networks, comparing with conventional LLM-based fine-tuning methods such as Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2021), as shown in Figure 2.
-We focus on two main issues when sequential full fine-tuning the pre-trained foundation models: (1) Catastrophic Forgetting (CF) and (2) parameter-efficient fine-tuning CL model.
-To deploy a practical model to deal with the two points, we suggest a new paradigm for Continual Learning (CL), named Well-initialized Lottery Ticket Hypotehesis:
+With the advent of large-scale pre-trained Transformers, i.e., ViT (Dosovitskiy et al., 2021), a new paradigm - (4) Prompt-based Continual Learning—has emerged.
+By freezing the backbone and learning lightweight prompts, these methods adapt the model to sequential tasks while preserving the strong representational priors of the pre-trained architecture.
+L2P (Wang et al., 2022d) first demonstrated that prompt retrieval can bridge prompting and continual representation learning.
+DualPrompt (Wang et al., 2022c) further decomposes prompts into task-general and task-specific components, improving adaptation and stability.
+Subsequent advances include DyTox (Douillard et al., 2022), S-Prompt (Wang et al., 2022a), CODA-P (Smith et al., 2023b), ConStruct-VL (Smith et al., 2023a), ST-Prompt (Pei et al., 2023), LGCL (Khan et al., 2023), ESN (Wang et al., 2022b), CPrompt (Gao et al., 2024), and NoRGa (Le et al., 2024).
+Despite their effectiveness, prompt-based CL methods exhibit several important limitations.
+First, performance depends heavily on the quality, granularity, and design of prompts, making the method sensitive to prompt initialization and capacity.
+Second, as the number of tasks grows, managing a large and diverse prompt pool becomes cumbersome.
+Third, prompt tuning inherently lacks the expressiveness of full fine-tuning-restricted prompts alone cannot capture the fine-grained task-specific variations needed for diverse and uncorrelated sequential tasks.
+While frozen pre-trained models offer strong global priors, they do not directly provide task-adaptive pathways, limiting the flexibility of prompt-only tuning.
+These limitations motivate the exploration of alternative fine-tuning paradigms that combine the stability of frozen backbones with the flexibility of full parameter adaptation.
+To address these challenges, we revisit the parameter-efficient fine-tuning landscape, especially in comparison to methods such as Adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2022), as illustrated in Figure˜2.
+We focus on two key desiderata in sequentially adapting pre-trained foundation models: (1) minimizing catastrophic forgetting and (2) ensuring parameter-efficiency during continual adaptation.
+Motivated by these goals, we introduce the following hypothesis:
 Well-initialized Lottery Ticket Hypothesis (WLTH).
-A well-initialized dense neural network contains globally minimal solutions that can retain the prior class knowledge while providing room to learn the new class knowledge through isolated fine-tuning of the networks or subnetworks.
-Leveraged by the WLTH, this work proposes a new Soft-TransFormer (Soft-TF) to address fine-tuning with minimal CF, as shown in Figure 1.
-In this work, we find task-specific soft-networks based on well-trained frozen transformer parameters that incrementally learn task-adaptive weights associated with sequential tasks.
-Our contributions can be summarized as follows:
-  • Inspired by Well-initialized Lottery Ticket Hypothesis (WLTH), we propose a novel continual learning method referred to as Soft-TransFormers (Soft-TF), which learns compact task-specific soft-networks from well pre-trained parameters for each task.
-  • Extensive experiments demonstrate that Soft-TF leads to better generalized parameter-efficient continual models than baselines such as DualPrompts, Adaptor, and LoRA, achieving state-of-the-art performances on Vision and Language Class-Incremental Learning (CIL) scenarios.
+A well-initialized dense neural network contains structured, globally stable solutions that preserve prior task knowledge, while also exposing local subspaces that can be selectively adapted for new tasks.
+These subspaces - corresponding to soft, real-valued subnetworks - enable isolated and parameter-efficient fine-tuning without overwriting the information encoded in the frozen backbone.
+Inspired by WLTH, we propose Soft-Transformer (Soft-TF), a new parameter-efficient continual learning framework that discovers task-specific soft subnetworks within well-trained Transformer backbones, as shown in Figure˜1.
+Soft-TF incrementally learns task-adaptive multiplicative masks that modulate key, query, value, and output projections in self-attention layers, enabling flexible and stable adaptation with minimal interference between tasks.
+Our main contributions are as follows:
+  • We propose Soft-Transformer (Soft-TF), a new continual learning paradigm inspired by the Well-initialized Lottery Ticket Hypothesis, which learns compact task-specific soft subnetworks from well pre-trained Transformers.
+  • We demonstrate that Soft-TF achieves state-of-the-art parameter-efficient continual learning performance, outperforming DualPrompt, NoRGa, Adapter, and LoRA across class-incremental learning scenarios.
 Refer to caption Refer to caption Refer to caption
-(a) Adapter (b) LoRA (c) Soft-TransFormer
-Figure 2: Comparisions of LLM-based fine-tuning methods (Adapters, LoRA) with Soft-TransFormer (Soft-TF): well-initialized Soft-TF $\bm{\theta}\odot\bm{m}_{t}$ is trained at $l_{e}$-th or only few attention layers while LLM fine-tuning methods such as Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2021) are trained at all attention layers.
+(a) Adapter (b) LoRA (c) our Soft-TransFormer
+Figure 2: Comparisions of LLM-based fine-tuning methods (Adapters, LoRA) with Soft-TransFormer (Soft-TF): well-initialized Soft-TF $\bm{\theta}\odot\bm{m}_{t}$ is trained at $l_{e}$-th or only few attention layers while LLM fine-tuning methods such as Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2022) are trained at all attention layers.
 Note that all pre-trained parameters $\bm{\theta}$ are fixed in fine-tuning t-th task parameter $\bm{m}_{t}$ in Continual Learning (CL) scenarios.
-2 Related Works
-Continual Learning (McCloskey and Cohen, 1989; Thrun, 1995; Kumar and Daume III, 2012; Li and Hoiem, 2016) is the challenge of learning a sequence of tasks continuously while utilizing and preserving previously learned knowledge to improve performance on new tasks.
-Four major approaches have been proposed to tackle the challenges of continual learning, such as catastrophic forgetting.
-One such approach is (1) Regularization-based approaches (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020; Titsias et al., 2020; Mirzadeh et al., 2021), which aim to reduce catastrophic forgetting by imposing regularization constraints that inhibit changes to the weights or nodes associated with past tasks.
-(2) Rehearsal-based approaches (Rebuffi et al., 2017; Chaudhry et al., 2019a, b; Saha et al., 2021; Deng et al., 2021; Sun et al., 2023; Sarfraz et al., 2023; Mai et al., 2021; Lin et al., 2023; Aljundi et al., 2019; Caccia et al., 2021; Chaudhry et al., 2019c; Liang and Li, 2024; Buzzega et al., 2020) store small data summaries to the past tasks and replay them during training to retain the acquired knowledge.
-Some approaches in this line of work (Shin et al., 2017; Aljundi et al., 2019) accommodate the generative model to construct the pseudo-rehearsals for previous tasks.
-(3) Architecture-based approaches (Mallya et al., 2018; Serrà et al., 2018; Li et al., 2019; Wortsman et al., 2020; Kang et al., 2022, 2023, 2024b, 2024a) use the additional capacity to expand (Xu and Zhu, 2018), dynamic representation (Yan et al., 2021; Singh et al., 2020) or isolate (Rusu et al., 2016) model parameters, preserving learned knowledge and preventing forgetting.
-Rehearsal and architecture-based methods have shown remarkable efficacy in suppressing catastrophic forgetting but require additional capacity for the task-adaptive parameters (Wortsman et al., 2020) or the replay buffers.
-Recently, (4) Prompt-based approaches, an emerging transfer learning technique, harnesses a fixed function of pre-trained Transformer models.
-This empowers the language model to receive additional instructions for enhancing its performance on downstream tasks.
-Two key works stand out in the prompt-based CL: L2P (Wang et al., 2022c), a seminal study bridging the gap between prompting and continual learning, and DualPrompt (Wang et al., 2022b), which proposes an innovative method for attaching complementary prompts to a fixed pre-trained backbone.
-Here, for the first time, we introduce a new approach to update the fixed pre-trained parameters through learnable sparse layered soft-networks under the convergence theory, maximumly enabling the acquisition of task-invariant and task-specific instructions with prompts.
-Prompt-based CL.
-With advances in Vision Transformers (Khan et al., 2022) and prompt-based fine-tuning in NLP (Li and Liang, 2021) and CL (Razdaibiedina et al., 2023), Wang et al.
-(2022c) have shown that interacting with a pre-trained model via prompt learning is a promising approach, L2P  (Wang et al., 2022c) for continual learning.
-Then, other advanced prompt-based CL models emerged such as DualPrompt (Wang et al., 2022b), DyTox (Douillard et al., 2022), S-Prompt (Wang et al., 2022a), CODA-P (Smith et al., 2023b), ConStruct-VL (Smith et al., 2023a), ST-Prompt (Pei et al., 2023), and LGCL (Khan et al., 2023).
-Recently, Prompt Gradient Projection (PGP) (Qiao et al., 2024), a small set of learnable orthogonal parameters, is appended to the input and enables quick adaptation of a frozen ImageNet pre-trained model to new streaming tasks.
-Their analysis shows that directly leveraging the pre-trained vision-language model without introducing any learnable parameters is a simple yet promising approach to continual learning.
-The PGP was also adopted to a joint vision-language model like CLIP (Radford et al., 2021) for continual learning, which presents multiple advantages - catering to practical scenarios without well-defined task identities and boundaries in class incremental learning scenarios.
-However, prompt-based models come with several disadvantages.
-Poorly trained prompts could lead to suboptimal performance or tend to be biased.
-Moreover, prompt tuning could not capture all nuances of uncorrelated sequential tasks.
-These disadvantages lead to exploring alternative methods like full fine-tuning or hybrid approaches for more robust and flexible prompt-based model performance.
-In this work, to alleviate these issues of prompt tuning, we investigate a fully fine-tuning of well-pre-trained transformers on training soft networks on various continual learning scenarios, comparing with LLM-based CL baselines such as MBPA++ (de Masson D’Autume et al., 2019), IDBR (Huang et al., 2021), and SLM (Bohao et al., 2024).
+2 Related Work
+Continual Learning.
+Continual Learning (CL) aims to enable models to accumulate knowledge over a sequence of tasks while preserving performance on previously learned ones (McCloskey and Cohen, 1989; Thrun, 1994).
+Early efforts focused on mitigating catastrophic forgetting, the tendency of neural networks to override earlier representations when trained on new data.
+Regularization-based methods address this by constraining parameter updates that would negatively impact past tasks, often by estimating parameter importance or imposing functional stability (Kirkpatrick et al., 2017; Chaudhry et al., 2020; Jung et al., 2020).
+Rehearsal-based approaches alleviate forgetting through explicit memory replay, either by storing exemplar sets or synthesizing pseudo-data from generative models (Rebuffi et al., 2017; Chaudhry et al., 2019a; Buzzega et al., 2020).
+Architecture-based strategies expand or isolate parts of the model so that new tasks use disjoint or sparsely overlapping parameter subsets, thereby preventing interference (Mallya et al., 2018; Serrà et al., 2018; Wortsman et al., 2020; Kang et al., 2022).
+Although these methods have proven effective, they often require additional storage, dynamic model growth, or explicit replay buffers, which can become impractical in long-horizon or resource-constrained CL scenarios.
+Parameter-efficient Adaptation.
+The shift from training models from scratch toward adapting large-scale pre-trained Transformers - such as ViT and CLIP (Dosovitskiy et al., 2021; Radford et al., 2021) - has significantly influenced recent CL research.
+Parameter-efficient fine-tuning (PEFT) methods, including adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2022), introduce compact trainable modules while keeping most backbone parameters frozen.
+These approaches provide stronger adaptation capacity than traditional CL techniques but still accumulate per-task modules or suffer from interference across tasks when task similarity is low.
+Prompt-based CL offers an even lighter adaptation mechanism: instead of modifying model weights, prompts act as learnable soft instructions prepended or inserted into the Transformer architecture.
+L2P (Wang et al., 2022d) initiated this direction by retrieving task-relevant prompts from a pool, while DualPrompt (Wang et al., 2022c) decomposed prompts into task-general and task-specific components to balance stability and plasticity.
+More recent advances aim to improve the expressiveness and robustness of prompt-based continual learning.
+CODA-Prompt (Smith et al., 2023b) utilizes decomposed attention to assemble input-conditioned prompts, mitigating task interference, while S-Prompt (Wang et al., 2022a) emphasizes simplicity and stability via domain-independent prompts.
+HiDe-Prompt (Wang et al., 2023) explores hierarchical and disentangled prompt representations to better capture task structures, and NoRGa (Le et al., 2024) introduces non-linear residual gating to enhance adaptability by connecting prompting with mixture-of-experts mechanisms.
+Collectively, these methods highlight the potential of prompt-based PEFT for continual learning, while also exposing inherent limitations in relying solely on prompt-level conditioning for scalable and expressive task adaptation.
+Limitations of Prompt-based CL.
+Despite their attractive efficiency, prompt-based methods face intrinsic constraints that limit their robustness in practical continual learning settings.
+Since prompts modify representations only implicitly through input conditioning, their expressiveness is fundamentally lower than module-level or weight-level adaptation, making them less suited for highly heterogeneous or weakly correlated task sequences.
+Moreover, prompt retrieval and initialization are sensitive to distribution shifts, and prompt pools may grow without bound as tasks accumulate, reducing scalability and increasing memory cost.
+Adapter- and LoRA-based PEFT methods provide stronger adaptation capacity but may introduce task interference or undesirable parameter growth.
+These challenges highlight the need for a principled and expressive mechanism that maintains the efficiency of PEFT while allowing direct modulation of the pre-trained backbone.
+Motivated by the Well-initialized Lottery Ticket Hypothesis (WLTH), we propose a Soft-Transformer (Soft-TF), which learns soft, real-valued subnetworks that multiplicatively modulate attention projections in a frozen Transformer.
+By selectively activating task-relevant subspaces within a well-initialized model, Soft-TF achieves stable, expressive, and parameter-efficient continual adaptation without relying solely on prompt proliferation.
+Figure 3: Soft-TransFormer (Soft-TF): At training time, the E-Prompt and the Soft-network are selected according to task identity, and the selected G-Prompt, E-Prompt, and the Soft-MSAs are trained together with a classifier.
+At test time, an input is transformed by a query function (Prompt ID) or task identifier to match the closest task key $\bm{k}_{t}$, E-prompt $\bm{e}_{t}$ and Soft-networks $\bm{m}_{t}^{\{K,Q,V\}}$.
 3 Prerequisites
-We start with conventional prompt-based continual learning methods using Vision Transformer (ViT) (Dosovitskiy et al., 2020) in Class Incremental Learning (CIL) scenarios.
+We start with conventional prompt-based continual learning methods using Vision Transformer (ViT) (Dosovitskiy et al., 2021) in Class Incremental Learning (CIL) scenarios.
 3.1 Problem Statement
@@
 3.2 Prompt-based Class Incremental Learning (CIL)
-A simple yet effective prompt-based (prompt-tuning) CIL model: Learning to Prompt (L2P) (Wang et al., 2022c) is first proposed.
-In this model, a prompt $p$, a tiny set of trainable tokens combined with image features, is fed into the Vision Transformer (ViT) to help the model resist forgetting.
-To select suitable prompts for task-specific training, L2P utilizes a prompt pool $P$ containing numerous prompt-key pairs, $\{p_{t},k_{t}\}_{t=1}^{\mathcal{T}}$, where ${p}_{t}\in\mathbb{R}^{1\times D}$ represents the $t$-th task prompt, $k_{t}$ represents the $t$-th coresponding task key, and $\mathcal{T}$ is the total number of prompt-key pairs.
-Building on L2P, DualPrompt (Wang et al., 2022b) divided the prompts into expert (E-) prompts and general (G-) prompts for distinct features learning.
-DualPrompt also replaced prompt-tuning with prefix-tuning, which was successfully proven in Language CIL scenarios.
-4 Transformer with Learnable Soft-networks
-In this section, we explain how Soft-TransFormers (Soft-TF) leverage learnable soft-networks to train sequential tasks while keeping the well-pretrained model parameters fixed.
-To introduce our novel Soft-TF and provide a clearer understanding, we draw on a partial explanation of DualPrompt.
+A simple yet effective prompt-based (prompt-tuning) CIL model: Learning to Prompt (L2P) (Wang et al., 2022d) is first proposed.
+In this model, a prompt $\bm{p}$, a tiny set of trainable tokens combined with image features, is fed into the Vision Transformer (ViT) to help the model resist forgetting.
+To select suitable prompts for task-specific training, L2P utilizes a prompt pool $\bm{P}$ containing numerous prompt-key pairs, $\{\bm{p}_{t},\bm{k}_{t}\}_{t=1}^{\mathcal{T}}$, where $\bm{p}_{t}\in\mathbb{R}^{1\times D}$ represents the $t$-th task prompt, $\bm{k}_{t}$ represents the $t$-th coresponding task key, and $\mathcal{T}$ is the total number of prompt-key pairs.
+Building on L2P, DualPrompt (Wang et al., 2022c) divided the prompts into expert (E-) prompts and general (G-) prompts for distinct features learning.
+DualPrompt also replaced prompt-tuning with prefix-tuning, which was successfully proven in CIL scenarios.
+4 Soft-TransFormer
+This section presents Soft-TransFormer (Soft-TF), a parameter-efficient continual learning framework that adapts a frozen pre-trained Transformer through continuous, learnable soft-subnetworks and structured multi-layer prompting.
+Soft-TF preserves the rich representational structure encoded in the pre-trained backbone while enabling flexible task-specific modulation with minimal computational overhead and significantly reduced forgetting.
 4.1 Soft-MSA Layers
-To address the task-specific fine-tuning of the pre-trained model, such as ViT, this work proposes a new Soft-TransFormer (Soft-TF), as illustrated in Figure 1.
-The proposed Soft-TF consists of a conventional neural network, like a multilayer transformer with multihead attention and forward networks.
-Using well-trained transformer parameters, we could discover task-specific soft-networks, as depicted in Figure 4.
-The Soft-TF incrementally learns model weights and task-adaptive soft-masks with well-pre-trained and soft-network parameters $\bm{m}$.
-The Soft-TF originates from learned parameters distributed with $\mu\approx 1.0$ & various variances, as stated in Figure 3: the histogram density plots compare the distribution of parameters for the attention layer L[12] in two contexts: the masking mechanism ($\bm{m}^{QKV}$) and the fine-tuned weighted product ($\bm{w}^{QKV}\odot\bm{m}^{QKV}$) for two tasks (Task 2 and Task 3).
-In Figure 3(a), the $\bm{m}^{QKV}$ parameters are concentrated in a narrow range centered around 1.
+Let $\bm{\theta}$ denote the parameters of a pre-trained Transformer, which remain fixed throughout continual learning.
+For each task $t$, as depicted in Figure˜3, Soft-TF introduces a set of real-valued masks
+$\bm{m}_{t}=\{\bm{m}_{t}^{Q},\,\bm{m}_{t}^{K},\,\bm{m}_{t}^{V},\,\bm{m}_{t}^{O}\},$
+which multiplicatively modulate the attention projections. Given hidden features $h$, we obtain:
+$\bm{Q}=h(\bm{\theta}^{Q}\odot\bm{m}_{t}^{Q}),\;\bm{K}=h(\bm{\theta}^{K}\odot\bm{m}_{t}^{K}),\;\bm{V}=h(\bm{\theta}^{V}\odot\bm{m}_{t}^{V}).$
+The output of the multihead attention layer is:
+$\mathrm{MSA}(h)=\mathrm{Attn}(\bm{Q}_{i},\bm{K}_{i},\bm{V}_{i})\Big(\bm{\theta}^{O}\odot\bm{m}_{t}^{O}),$ (1)
+where $\mathrm{Attention}(\cdot)$ denotes standard scaled dot-product attention.
+Intuition.
+The masks are initialized at $\mathbf{1}$, ensuring that the initial computation matches the pre-trained model exactly.
+During optimization, the masks induce task-specific, continuous subnetwork selection within each attention projection.
+This allows Soft-TF to adjust representational pathways while keeping the global structure of the backbone fixed, thereby constraining parameter drift and enabling stable adaptation across tasks.
+4.2 Prompts Coupled with Soft-Subnetwork
+Soft-TF incorporates both general prompts (shared across tasks) and expert prompts (task-specific) into a multi-layer prompting mechanism.
+Let $\bm{g}=\{\bm{g}^{l}\}_{l=l_{g}^{\text{start}}}^{l_{g}^{\text{end}}}$, $\bm{e}_{t}=\{\bm{e}_{t}^{l}\}_{l=l_{e}^{\text{start}}}^{l_{e}^{\text{end}}}$ denote the layerwise G-prompts and E-prompts, respectively, with the constraint $l_{g}^{\text{start}},l_{g}^{\text{end}}]\cap[l_{e}^{\text{start}},l_{e}^{\text{end}}]=\emptyset.$
+Given an input sequence $\bm{x}$, prompts are concatenated as
+$\bm{z}=[\,\bm{p};\,\bm{x}\,],\quad\bm{p}\in\{\bm{g},\bm{e}_{t}\}.$
+Soft-MSA then processes $\bm{z}$ via masked projections:
+$\bm{q}=(\bm{\theta}^{Q}\odot\bm{m}_{t}^{Q})\bm{z},\qquad\bm{k}=(\bm{\theta}^{K}\odot\bm{m}_{t}^{K})\bm{z},$
+yielding the attention matrix
+$\bm{a}=\mathrm{softmax}\left(\frac{\bm{q}\bm{k}^{\top}}{\sqrt{D/s}}\right).$ (2)
+G-prompts impart globally relevant cues, while E-prompts and soft-masks enable fine-grained, task-specific adjustments.
+This combination yields strong task separation without requiring explicit task identifiers during evaluation.
+4.3 Optimization over Frozen Backbones
+For task $t$, the predictive model is
+$f_{t}(\bm{x})=f_{\phi}\big(f_{\bm{\theta},\bm{g},\bm{e}_{t},\bm{m}_{t}}(x)\big),$
+where $\phi$ is a lightweight task classifier. Training optimizes:
+$\min_{\bm{g},\,\bm{e}_{t},\,\bm{m}_{t},\,\bm{k}_{t},\,\phi}\;\mathcal{L}_{\mathrm{CE}}\big(f_{\phi}(f_{\bm{\theta},\bm{g},\bm{e}_{t},\bm{m}_{t}}(\bm{x})),y\big)+\lambda\,\gamma\big(q(\bm{x}),\bm{k}_{t}\big),$ (3)
+where $\bm{k}_{t}$ is a task-level key vector and $\gamma$ is a cosine-similarity regularizer, defined as $1-\cos(\cdot,\cdot)$, which facilitates task inference at test time.
+Only the prompts, soft-masks, and classifier parameters are updated; the backbone remains frozen, yielding low memory cost and consistent forward-time efficiency.
+The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm˜1 and Algorithm˜2.
+Algorithm 1 Training Procedure of DualPrompt-Soft-TF
+1:  Input: Pre-trained transformer backbone $f$, classifier $f_{\phi}$;
+2:   number of tasks $\mathcal{T}$; training sets $\{\{(\bm{x}_{i,t},y_{i,t})\}_{i=1}^{n_{t}}\}_{t=1}^{\mathcal{T}}$;
+3:   G-Prompt $\bm{g}$, E-Prompts $\bm{E}=\{\bm{e}_{t}\}_{t=1}^{\mathcal{T}}$;
+4:   task keys $\bm{K}=\{\bm{k}_{t}\}_{t=1}^{\mathcal{T}}$;
+5:   soft-networks $\bm{M}=\{\bm{m}_{t}\}_{t=1}^{\mathcal{T}}$;
+6:   prompt insertion layers $l^{start}_{g},l^{end}_{g}]$, $l^{start}_{e},l^{end}_{e}]$;
+7:   prompting function $f_{\bm{\theta}\odot\bm{m}}$; epochs $\mathcal{K}_{t}$ for task $t$.
+8:  Initialize: $\phi$, $\bm{g}$, $\bm{E}$, $\bm{M}$, $\bm{K}$
+9:  for task $t=1,\dots,\mathcal{T}$ do
+10:  Select task-specific $\bm{e}_{t}$, soft-network $\bm{m}_{t}$, and key $\bm{k}_{t}$
+11:  Construct prompted backbone $f_{\bm{g},\bm{e}_{t},\bm{m}_{t}}$ by attaching:
+12:   G-Prompt to layers $l^{start}_{g},l^{end}_{g}]$,
+13:   E-Prompt to soft-MSA layers $l^{start}_{e},l^{end}_{e}]$ via $f_{\bm{\theta}\odot\bm{m}_{t}}$
+14:  for epoch $s=1,\dots,\mathcal{K}_{t}$ do
+15:   Sample mini-batch $B=\{(\bm{x}_{i,t},y_{i,t})\}_{i=1}^{l}$
+16:   for each $(\bm{x},y)$ in $B$ do
+17:    Compute prompted representation:
+18:     $\bm{h}=f_{\bm{g},\bm{e}_{t},\bm{m}_{t}}(\bm{x})$
+19:    Compute prediction $\hat{y}=f_{\phi}(\bm{h})$
+20:    Compute per-sample loss:
+21:     $\mathcal{L}_{x}=\mathcal{L}_{CE}(\hat{y},y)+\lambda\,\gamma(q(\bm{x}),\bm{k}_{t})$
+22:   end for
+23:   Update $\phi$, $\bm{g}$, $\bm{E}$, $\bm{M}$, $\bm{K}$ via backpropagation
+24:  end for
+25:  end for
+4.4 Stability and Bounded Parameter Drift
+Define the effective parameters for task $t$ as
+$\bm{w}=\bm{\theta}\odot\bm{m}_{t},$
+so that $\bm{w}^{(1)}=\bm{\theta}$ under mask initialization $\bm{m}_{t}^{(1)}=\mathbf{1}$. A single gradient update yields:
+$\bm{w}^{(k+1)}=\bm{w}^{(k)}-\eta\big(\bm{\theta}\odot\nabla_{\bm{m}}\mathcal{L}\big).$
+Thus, the optimization path is governed by a rescaled gradient whose magnitude is bounded by the scale of $\bm{\theta}$.
+Under standard convex–Lipschitz conditions, we obtain the cumulative suboptimality bound:
+$\begin{aligned} \sum_{k=1}^{T}&\langle\bm{w}^{(k)}-\bm{w}^{\star},\,\bm{\theta}\odot\nabla_{\bm{m}}\mathcal{L}\rangle\leq\\ &\mathcal{O}\!\left(\frac{\|\bm{w}^{(1)}-\bm{w}^{\star}\|^{2}}{\eta}+\eta\sum_{k=1}^{T}\|\bm{\theta}\odot\nabla_{\bm{m}}\mathcal{L}\|^{2}\right).\end{aligned}$
+Table 1: Performances of Vision Class Incremental Learning in terms of Accuracy, Forgetting, Trainable parameters, and Time complexity on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R.
+Note that Soft-TF performed without FeedForward ($F$) networks.
+$\ast$ denotes our reproduced results and L[10,11,12] represents Soft-TransFormer of 10, 11, 12 Layers.
+ViT-B/16 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
+Method #Tr.Params. Task ID ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec]
+L2P 0.03M Prompt ID 83.77 / 6.63 12.00K / 75 71.29 / 13.96 11.50K / 76 60.44 / 9.00 12.80K / 46
++ PGP 0.03M Prompt ID 84.34 / 5.59 12.30K / 75 76.12 / 13.26 12.70K / 76 61.40 / 8.03 13.07K / 46
++ Soft-TF-L[3,4,5] 6.93M Prompt ID 86.26 / 4.79 12.85K / 78 76.17 / 15.77 13.98K /100 69.80 / 5.13 14.23K / 49
+DualPrompt 0.03M Prompt ID 86.50 / 5.77 12.12K / 76 82.98 / 8.20 11.60K / 78 68.13 / 4.46 13.10K / 47
++ PGP 0.03M Prompt ID 86.92 / 5.35 12.21K / 76 83.74 / 7.91 13.12K / 78 63.34 / 4.53 13.33K / 47
++ Adapter 6.93M Prompt ID 86.51 / 4.75 14.98K / 90 84.48 / 5.81 16.13K / 109 70.56 / 4.71 15.65K / 54
++ LoRA 6.93M Prompt ID 82.00 / 4.33 13.24K / 79 92.14 / 2.02 15.89K / 105 43.51 / 13.2 15.09K / 53
++ Soft-TF-L[10,11,12] 6.93M Prompt ID 92.35 / 2.98 13.87K / 80 97.40 / 0.57 15.60K / 104 76.62 / 5.30 15.35K / 52
+NoRGa 0.03M Prompt ID 94.48 / 1.44 12.12K / 76 97.39 / 0.58 11.60K / 78 75.40 / 4.59 13.10K / 47
++ Soft-TF-L[10,11,12] (SOTA) 6.93M Prompt ID 98.25 / 0.31 13.87K / 80 98.30 / 0.30 15.60K / 104 91.94 / 0.34 15.35K / 52
+Joint of DualPrompt 0.13M - 90.85 - 90.85 - 79.13 -
+Joint of DualPrompt+Soft-TF 6.93M - 95.90 - 95.90 - 80.21 -
+Joint of NoRGa+Soft-TF 6.93M - 98.20 - 98.20 - 90.73 -
+Implications for continual learning.
+The initial distance $\|\bm{w}^{(1)}-\bm{w}^{\ast}\|$ is naturally small across most tasks because the pre-trained backbone already provides strong feature representations.
+Consequently, by standard convergence guarantees of first-order optimization for Lipschitz-continuous objectives without assuming strong convexity, the suboptimality decreases at a rate of $\mathcal{O}(1/\sqrt{T})$.
+More importantly, since optimization occurs only in the low-dimensional soft-parameter space $(\bm{m}_{t}^{Q},\bm{m}_{t}^{K},\bm{m}_{t}^{V})$ while $\bm{\theta}$ remains fixed, the entire trajectory $\bm{w}^{(k)}$} stays confined to a narrow region around the pretrained solution manifold.
+This imposes an inherent bound on parameter drift and explains why Soft-TF reliably preserves global knowledge encoded in $\bm{\theta}$ while still enabling targeted adaptation to each new task.
+Taken together, these theoretical results (see Section˜A.1) provide a principled foundation for our design choice: initializing masks at $\mathbf{1}$ and fine-tuning only $\bm{m}_{t}$ over a frozen backbone yields stable optimization, bounded deviation from pre-trained representations, and substantially reduced catastrophic forgetting in continual learning.
+5 Experiments
+We validate our Soft-TF on several benchmark datasets against continuous learning baselines in Class-Incremental Learning (CIL).
+5.1 Experimental Settings
+Datasets.
+We evaluate our method mainly on 1) 10/20-Split-CIFAR100 (Krizhevsky et al., 2009), constructed by splitting the 100 classes into 10 tasks/20 tasks.
+2) 10-Split-TinyImageNet (Abai and Rajmalwar, 2019), constructed by splitting the 200 classes into 10 tasks.
+3) 10-Split-ImageNet-R (Hendrycks et al., 2021), constructed by splitting the 200 classes into 10 tasks.
+To show our effectiveness, we additionally compare our method with the baselines on 5-Split-CUB200.
+Baselines.
+To validate the powerfulness of our method, we compare our results with various CIL baselines including, L2P (Wang et al., 2022d), DualPrompt (Wang et al., 2022c), DualPrompt-PGP (Qiao et al., 2024), and NoRGa (Le et al., 2024).
+At inference time, the system determines the task identity for an arbitrary test sample $\bm{x}$ to select the appropriate task-specific nuances and ultimately achieve full fine-tuning performance.
+This is accomplished by using the Prompt ID mechanism, which simply chooses the best-matched task index $t$ by finding the minimum value of a function $\gamma$ comparing the query representation of the sample, $q(\bm{x})$, against all learned task keys $\bm{k}_{t}$, expressed as $\text{argmin}_{t}\gamma(q(\bm{x}),\bm{k}_{t})$.
+Table 2: Performances of DualPrompt+Soft-TF in terms of ViTs pretained-dataset (ImageNet-21K) on 10-Split-ImageNet-R.
+DualPrompt 10-Split-ImageNet-R
++Soft-TF #Tr.Prams ACC($\uparrow$) Forget($\downarrow$) Tr./Test[sec]
+ViT-Ti/16 (5.7M) 0.41M 49.34 9.02 15.10K / 50
+ViT-S/16 (22M) 1.71M 65.12 6.67 15.22K / 51
+ViT-B/16 (86M) 6.93M 76.62 5.30 15.35K / 52
+ViT-L/16 (307M) 12.50M 78.86 4.46 28.57K / 99
+Refer to caption Refer to caption
+(a) Comparisons with baselines (b) Layer-wise Performances
+Figure 4: Layer-wise(L[$\ast$]) Performances of DualPrompt+Soft-TF on 10-Split-CIFAR100. Note that L[10,11,12] denotes Soft-Transformer of 10, 11, 12 Layers.
+5.2 Performances
+Performances of Soft-TF.
+We compare our Soft-TF with state-of-the-art CIL baselines, as shown in Table˜1.
+Soft-TF consistently outperforms prompt-based and parameter-efficient fine-tuning methods—including L2P, DualPrompt, NoRGa, Adapter, LoRA, and their corresponding upper-bounds—across accuracy, forgetting, and efficiency metrics.
+In particular, Soft-TF exhibits substantial improvements over NoRGa, a strong prompt-based baseline that mitigates task interference, yet remains limited by its reliance on global or head-level prompt modulation.
+By contrast, Soft-TF directly modulates attention projection matrices through task-specific soft subnetworks, enabling fine-grained and continuous adaptation within a frozen Transformer backbone, which results in higher accuracy and significantly reduced forgetting across all benchmarks.
+Despite using comparable trainable parameter budgets, Soft-TF achieves these gains without additional training or inference cost.
+Remarkably, Soft-TF even surpasses the Upper-Bound (Joint) of DualPrompt using prompt baselines: DualPrompt and NoRGa on all evaluated benchmarks, as it explicitly identifies and optimizes distinct task-specific soft subnetworks for each incremental task, achieving near-optimal per-task specialization, whereas joint training is constrained to a single shared solution that must compromise across tasks.
+Comparisions with baselines.
+Figure˜4 of (a) shows that Soft-TF consistently outperforms established parameter-efficient tuning baselines - including PGP, Adapter, and LoRA - across all sessions in 10-Split CIFAR-100, demonstrating superior resistance to task interference and better retention of previously learned knowledge.
+As shown in Figure˜4 of (b) the layer-wise analysis, even when fine-tuning only selected layers (e.g., layers 10, 11, 12), Soft-TF exhibits a range of performance behaviors depending on the chosen layer configuration, yet it still maintains a clear advantage over LoRA in every setting.
+This indicates that Soft-TF not only provides stronger overall continual learning stability but also preserves its effectiveness under various layer-wise adaptation strategies, highlighting its robustness and versatility.
+5.3 Variants of ViTs.
+Across different ViT backbones, Soft-TF exhibits consistent advantages in both parameter efficiency and continual learning stability on 10-Split-ImageNet-R, as shown in Table˜2.
+The lightweight ViT-Ti/16 uses only 0.41M trainable parameters but yields lower accuracy (49.34%) and higher forgetting (9.02), whereas scaling up to ViT-S/16, ViT-B/16, and ViT-L/16 increases trainable parameters to 1.71M, 6.93M, and 12.50M, respectively, resulting in markedly improved performance.
+In particular, Soft-TF with ViT-L/16 achieves the best overall results - 78.86% accuracy and only 4.46 forgetting - while maintaining comparable training and test times ($\approx$28K/99 sec).
+Figure˜5 shows that larger backbones maintain higher average accuracy across all 10 sessions, with ViT-L/16 degrading the slowest and sustaining a clear performance margin over ViT-B/16, ViT-S/16, and ViT-Ti/16.
+Overall performance differences are further influenced by each backbone’s representational capacity and initialization quality, which shape how effectively Soft-TF can adapt while resisting catastrophic forgetting.
+Moreover, while increasing the number of trainable parameters generally leads to better continual learning performance, this improvement naturally comes with higher computational cost, reflecting the typical trade-off between model capacity and efficiency.
+Figure 5: Performances of DualPrompt+Soft-TF: in terms of ViT models on 10-Split-ImageNet-R. The performance of Soft-TF-ViT-B/16 is comparable to that of Soft-TF-ViT-L/16.
+5.4 Well-initialized Soft-TF
+Random initialization.
+Random initialization of Soft-Transformer’s weights plays a critical role when leveraging well-pretrained models like Vision Transformers (ViTs).
+The optimal training point is the parameters of a well-pretrained model.
+Among the initialization methods, Uniform initialization for Soft-TransFormer satisfies this requirement effectively.
+To validate these claims, we analyze the impact of common random initialization methods, including Xavier, Kaiming, Normal, and Uniform Initialization, as shown in Table˜3.
+The results demonstrate that the same well-initialization point leads to independent optimal task performance.
+Furthermore, this ablation study strengthens our Soft-TF with state-of-the-art-perfomances inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH).
+Well-initialized LTH (WLTH).
+Based on Table˜4, we evaluated our Soft-TF on the 10-Split-CIFAR100 and 5-Split-CUB200 datasets using three distinct pre-trained ViT models (ImageNet-21K, SAM, and DINO) to demonstrate its efficacy, including on Well-initialized Lottery Ticket Hypothesis (WLTH) backbones.
+For the 10-Split-CIFAR100 benchmark, the ImageNet-21K backbone achieved the highest accuracy (92.35%) and the lowest forgetting rate (2.98%), closely followed by the SAM backbone (92.30% ACC, 2.99% Forget).
+Similarly, for the 5-Split-CUB200 benchmark, the ImageNet-21K backbone led with 76.17% ACC and 9.04% Forget, which was also nearly matched by the SAM backbone (76.11% ACC, 9.01% Forget).
+In contrast, the DINO backbone consistently showed the lowest performance across both datasets.
+These trends are consistent with our theoretical analysis, which predicts faster convergence and bounded parameter drift when optimization is initialized near a well-pretrained solution and confined to a low-dimensional soft-parameter space.
+These results confirm the effectiveness of Soft-TF and suggest that using well-initialized weights provides superior generalization and performance in continual learning settings.
+Table 3: Random initialized Performances of DualPrompt+Soft-TF on 10-Split-CIFAR100. Note DualPrompt’s parameters are pretrained on ImageNet-21K.
+10-Split-CIFAR100
+Random Init. ACC($\uparrow$) Forget($\downarrow$)
+Xavier 90.59 3.85
+Kaiming 90.72 3.63
+Normal 90.45 3.78
+Uniform(1.0, 1.0) 92.35 2.98
+Table 4: Performances of Soft-TF in terms of Pretained-dataset (ImageNet-21K, SAM, DINO) on 10-Split-CIFAR100 and 5-Split-CUB200.
+DualPrompt 10-Split-CIFAR100 5-Split-CUB200
++Soft-TF ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
+ImageNet-21K 92.35 2.98 76.17 9.04
+SAM 92.30 2.99 76.11 9.01
+DINO 84.50 12.27 69.79 10.93
+5.5 Soft-MSA Layers
+The Soft-TF originates from learned parameters distributed with $\mu\approx 1.0$ & various variances, as stated in Figure˜6: the histogram density plots compare the distribution of parameters for the attention layer L[12] in two contexts: the masking mechanism ($\bm{m}^{QKV}$) and the fine-tuned weighted product ($\bm{\theta}^{QKV}\odot\bm{m}^{QKV}$) for two tasks (Task 2 and Task 3).
+In Figure˜6(a), the $\bm{m}^{QKV}$ parameters are concentrated in a narrow range centered around 1.
 There is a slight variation between Task 2 and Task 3.
@@
 $\bm{m}^{QKV}$ acts as a soft gating mechanism, controlling which pre-trained parameters are retained or modified during fine-tuning.
-In LABEL:{fig:soft_weight_cifar10}(b), the distribution of pre-trained weight ($\bm{w}^{QKV}$) is centered around 0, with a Gaussian-like spread.
-The distribution of soft-masked fine-tuned weight ($\bm{w}^{QKV}\odot\bm{m}^{QKV})$ remain similar to that of $\bm{w}^{QKV}$ but slightly narrower.
+In LABEL:{fig:soft_weight_cifar10}(b), the distribution of pre-trained weight ($\bm{\theta}^{QKV}$) is centered around 0, with a Gaussian-like spread.
+The distribution of soft-masked fine-tuned weight ($\bm{\theta}^{QKV}\odot\bm{m}^{QKV})$ remain similar to that of $\bm{\theta}^{QKV}$ but slightly narrower.
 The overlap between the two distributions indicates that the fine-tuning process maintains the general structure of the pre-trained weights while introducing task-specific modifications.
@@
 Refer to caption Refer to caption
-(a) $\bm{m}^{QKV}$ (b) $\bm{w}^{QKV}$ v.s. $\bm{w}^{QKV}\odot\bm{m}^{QKV}$
+(a) $\bm{m}^{QKV}$ (b) $\bm{\theta}^{QKV}$ v.s. $\bm{\theta}^{QKV}\odot\bm{m}^{QKV}$
 Figure 6: Attention Layer-(L[12]) Histogram Density Estimates of Pre-trained weight ($\bm{\theta}$) and Soft-TF ($\bm{m}$) on 10-Split-CIFAR100.
-Given a pre-trained parameter $\bm{\theta}$ and learnable soft-parameters $\bm{m}$, Soft-ViT is represented as $f_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$, consisting of $N$ consecutive soft-MSA layers.
-We extend the notation by denoting the input embedding feature of the $l_{\ast}$-th learnable soft-MSA layer as $\bm{h}^{(l_{\ast})}$, where $l_{\ast}=1,2,\dots,N$, and $l_{\ast}$ can refer to either the G-Prompt layer $l_{g}$ or the E-Prompt layer $l_{e}$.
-Note that while the pre-trained parameters $\bm{\theta}$ remain fixed, the soft-parameters $\bm{m}$ are updated to provide task-specific solutions.
-G-prompt.
-$\bm{g}\in\mathbb{R}^{L_{g}\times D}$ with sequence length $L_{g}$ and embedding dimension $D$, is a shared parameter for all tasks.
-G-Prompt is attached to the $l_{g}$-th MSA layer to transform $\bm{h}^{(l_{g})}$ via a prompting function as follows:
-$\bm{h}_{g}^{(l_{g})}=f^{prompt}_{\bm{\theta}}(\bm{g},\bm{h}^{(l_{g})}),$
-where $f^{prompt}_{\bm{\theta}}$ defines the approach for attaching the prompt to the hidden embeddings.
-E-prompt & Soft-networks.
-$\bm{e}=\{\bm{e}_{t}\}^{\mathcal{T}}_{t=1}$ is a set of task-dependent parameters, where $\bm{e}_{t}\in\mathbb{R}^{L_{e}\times D}$ has as sequence length of $L_{e}$ and the same embedding dimension $D$ as the G-prompt, and $\mathcal{T}$ is the total number of tasks.
-Unlike the shared G-prompt, each $\bm{e}_{t}$ is associated with a task-specific key $\bm{k}_{t}\in\mathbb{R}^{D}$, which is also a learnable parameter aimed at capturing representative features of a task.
-For an input example from the $t$-th task, to attach E-prompt to the $l_{e}$-th soft-MSA layer, we apply the prompting function in a similar way:
-$\bm{h}_{e}^{(l_{e})}=f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}(\bm{e}_{t},\bm{h}^{(l_{e})}).$
-Figure 4: Soft-TransFormers (Soft-TF): At training time, the E-Prompt and the Soft-network are selected according to task identity, and the selected G-Prompt, E-Prompt, and the Soft-networks (Soft-Attention and Feed Forwards) are trained together with a classifier.
-At test time, an input is transformed by a query function (Prompt ID) or task identifier (Gradient ID) to match the closest task key $\bm{k}_{t}$, E-prompt $\bm{e}_{t}$ and Soft-networks $\bm{m}_{t}^{\{K,Q,V\}}$.
-Note task identifier is depicted in Section 5.
-4.2 Prompts with Learnable Subnetworks
-G- and E-prompts, along with learnable soft-networks, encode specific types of instructions during training with the backbone and work together to guide the model’s predictions during inference.
-We have demonstrated the method for attaching prompts and learnable soft-networks to a single soft-MSA layer.
-Similarly to the approach taken in DualPrompt (Wang et al., 2022b), we also investigate layers of E-prompts with learnable soft-networks $\bm{m}$, while utilizing the layers designated for G-prompts.
-Layers of G- and E-Prompts.
-We use the multilayered extension of both types of prompts: $\bm{g}=\{\bm{g}^{(l_{g})}\}_{l_{g}=start_{g}}^{end_{g}}$, where $\bm{g}^{(l_{g})}\in\mathbb{R}^{L_{g}\times D}$ represents the G-prompt attached to the $l_{g}$-th MSA layer.
-Similarly, we define $\bm{e}_{t}=\{\bm{e}_{t}^{(l_{e})}\}_{l_{e}=start_{e}}^{end_{e}}$ for the $l_{e}$-th conventional MSA layer.
-In this configuration, the G-prompt $\bm{g}^{(l_{g})}$ is attached from the $start_{g}$-th to the $end_{g}$-th conventional MSA layers, and the E-prompt $\bm{e}_{t}^{(l_{e})}$ is attached to the $\left[start_{e},end_{e}\right]$-th soft-MSA layers, ensuring that there is no overlap between them.
-In our experiments, we follow the $l_{g}\notin\left[start_{e},end_{e}\right]$ ($l_{g}=[1,2]$) settings used in DualPrompt and empirically search for the optimal $\left[start_{e},end_{e}\right]$ layers for the learnable subnetworks through ablation studies.
-Learnable Soft-networks.
-The prompting function $f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}$ determines how prompts ($\bm{p}$) are combined with fine-tuned soft ($\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}$) embedding features.
-From another perspective, $f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}$ directly influences the interaction between high-level instructions in the prompts and low-level representations.
-Therefore, we believe that a well-designed prompting function, along with task-specific parameters, is crucial for optimizing overall continual learning performance.
-Specifically, applying a prompting and fine-tuning function $f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}$ can be seen as modifying the inputs to the soft-MSA layers.
-Let the input to the soft-MSA layer be $\bm{h}\in\mathbb{R}^{L\times D}$, and denote the input query, key, and values for the soft-MSA layer as $\bm{h}_{Q}$, $\bm{h}_{K}$, and $\bm{h}_{V}$, respectively.
-A soft-MSA layer is defined by the following equation:
-$\text{MSA}(\bm{h}_{Q},\bm{h}_{K},\bm{h}_{V})=[\bm{h}_{1},\cdots,\bm{h}_{i},\cdots,\bm{h}_{n}]\bm{w}^{O}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{O}}$
-where $\bm{h}_{i}=\text{Attention}(\bm{h}_{Q}(\bm{w}^{Q}_{i}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{Q}}),\bm{h}_{K}(\bm{w}^{K}_{i}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{K}}),\bm{h}_{V}(\bm{w}^{V}_{i}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{V}}))$; $\bm{w}_{i}^{O},\bm{w}_{i}^{Q},\bm{w}_{i}^{K}$, and $\bm{w}_{i}^{V}$ are fixed projection matrices while ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{O}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{Q}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{K}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{V}$} are learnable parameters.
-$s$ is the number of heads.
-In ViT, $\bm{h}_{Q}=\bm{h}_{K}=\bm{h}_{V}$.
-Here, we define a unified prompt parameter with a sequence length of $L_{p}$, such as $\bm{p}\in\mathbb{R}^{L_{p}\times D}$ for a single-layered G- or E-prompt.
-4.3 Fine-tuning on Well-initialized Parameters
-In this framework, we concatenate the prompts $\bm{p}_{t}$ and the embedding sequence $\bm{x}_{t}$, i.e., inputs from $t$-th task, along the embedding dimension: $\bm{z}_{t}=[\bm{p}_{t};\bm{x}_{t}]$.
-With the weights of $\bm{w}^{Q}\odot\bm{m}^{Q},\bm{w}^{K}\odot\bm{m}^{K},\bm{w}^{V}\odot\bm{m}^{V}$, the soft-transformer takes query ($\bm{q}_{t}=(\bm{w}^{Q}\odot\bm{m}^{Q})\bm{z}_{t}$) and key ($\bm{k}_{t}=(\bm{w}^{K}\odot\bm{m}^{K})\bm{z}_{t}$) as input of the soft-MSA layer.
-The soft-attention matrix is then given by:
-$\bm{a}_{t}=\text{softmax}\left(\frac{\bm{q}_{t}\bm{k}_{t}^{T}}{\sqrt{D/n}}\right)$
-where we focus on $\bm{q}_{t}\bm{k}_{t}^{T}=(\bm{w}^{Q}\odot\bm{m}^{Q})\bm{z}_{t}$ and $\bm{z}_{t}^{T}(\bm{w}^{K}\odot\bm{m}^{K})^{T}$. First, the trainable prompt parameters can be denoted as:
-$\bm{z}_{t}\cdot\bm{z}_{t}^{T}=\begin{bmatrix}\bm{p}_{t}\\ \bm{x}_{t}\end{bmatrix}\begin{bmatrix}\bm{p}_{t}\;\bm{x}_{t}\end{bmatrix}=\begin{bmatrix}\bm{p}_{t}\bm{p}^{T}_{t}\;\bm{p}_{t}\bm{x}_{t}^{T}\\ \bm{x}_{t}\bm{p}_{t}^{T}\;\bm{x}_{t}\bm{x}_{t}^{t}\end{bmatrix}$
-Second, the trainable soft-attention layer’s parameters with $\bm{m}^{Q}$ and $\bm{m}^{K}$ are as follows:
-$\left\{\begin{matrix}\bm{m}^{Q}\cdot\bm{p}_{t}\bm{p}_{t}^{T}\cdot(\bm{m}^{K})^{T},\\ \bm{m}^{Q}\cdot\bm{x}_{t}\bm{p}_{t}^{T}\cdot(\bm{m}^{K})^{T},\\ \bm{m}^{Q}\cdot\bm{p}_{t}\bm{x}_{t}^{T}\cdot(\bm{m}^{K})^{T},\\ \bm{m}^{Q}\cdot\bm{x}_{t}\bm{x}_{t}^{T}\cdot(\bm{m}^{K})^{T}.\\ \end{matrix}\right\}$
-where $\bm{w}^{Q}$ and $\bm{w}^{K}$ are frozen and unchanged during training and test.
-The trainable soft-attention layer’s parameters $\bm{m}^{\ast}$ are used to analyze the convergence rate of the Soft-Transformer (see Section A.1 stated in detail).
-4.4 Optimization of Soft-Transformers
-The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm 1 and Algorithm 2.
-We denote the architecture with attached prompts as $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$.
-The input $\bm{x}$ of the $t$-th task is transformed using $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$ and then passed to the classification head $f\phi$, parameterized by $\phi$, for prediction.
-Finally, we train the two prompts, the task keys, the soft-attention parameters, and the newly-initialized classification head in an end-to-end manner:
-$\min_{\bm{g},\bm{e}_{t},\bm{m}_{t},\bm{k}_{t},\bm{\phi}}\mathcal{L}(f_{\phi}(f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}(\bm{x})),y)+\lambda\mathcal{L}_{\text{match}}(\bm{x},\bm{k}_{t}),$
-Here, $\mathcal{L}$ represents the cross-entropy loss, and $\mathcal{L}_{\text{match}}(\bm{x},\bm{k}_{t})=\gamma(\bm{q}(\bm{x}),\bm{k}_{t})$ denotes the matching loss, where $\bm{q}(\bm{x})=f(\bm{x})[0]$ corresponds to the feature vector associated with the [class] token (Dosovitskiy et al., 2020; Wang et al., 2022b), and $\gamma$ is the cosine similarity.
-The scalar $\lambda$ serves as a balancing factor between the losses; here, we follow the same DualPrompt setting as a baseline.
-Table 1: Performances of Vision Class Incremental Learning (CIL, ViTs) in terms of Accuracy, Forgetting, Trainable parameters, and Time complexity on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R.
-Note that Soft-TF performed without FeedForward ($F$) networks.
-$\ast$ denotes our reproduced results.
-ViT-B/12 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
-Method #Tr.Params. Task ID ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec]
-L2P 0.03M Prompt ID 83.77 / 6.63 12.00K / 75 71.29 / 13.96 11.50K / 76 60.44 / 9.00 12.80K / 46
- + PGP 0.03M Prompt ID 84.34 / 5.59 12.30K / 75 76.12 / 13.26 12.70K / 76 61.40 / 8.03 13.07K / 46
- + Soft-TF-L[3,4,5] 6.93M Prompt ID 86.26 / 4.79 12.85K / 78 76.17 / 15.77 13.98K /100 69.80 / 5.13 14.23K / 49
- + Soft-TF-L[10,11,12] 6.93M Gradient ID 86.46 / 4.87 12.87K / 128 77.67 / 15.84 13.95K / 160 69.56 / 5.28 14.23K / 79
-DualPrompt 0.03M Prompt ID 86.50 / 5.77 12.12K / 76 82.98 / 8.20 11.60K / 78 68.13 / 4.46 13.10K / 47
- + PGP 0.03M Prompt ID 86.92 / 5.35 12.21K / 76 83.74 / 7.91 13.12K / 78 63.34 / 4.53 13.33K / 47
- + Adapter-L[all], r=75 6.93M Prompt ID 86.51 / 4.75 14.98K / 90 84.48 / 5.81 16.13K / 109 70.56 / 4.71 15.65K / 54
- + LoRA-L[10,11,12], r=500 6.93M Prompt ID 82.00 / 4.33 13.24K / 79 92.14 / 2.02 15.89K / 105 43.51 / 13.21 15.09K / 53
- + Soft-TF-L[10,11,12] 6.93M Prompt ID 92.35 / 2.98 13.87K / 80 97.40 / 0.57 15.60K / 104 76.62 / 5.30 15.35K / 52
- + Soft-TF-L[10,11,12] 6.93M Gradient ID 98.05 / 0.25 13.71K / 130 98.96 / 0.23 15.51K / 163 83.70 / 0.53 15.08K / 82
-5 Experiments
-We validate our Soft-TF on several benchmark datasets against continuous learning baselines in (1) Class-Incremental Learning (CIL) based on ViT models  [^1] https://github.com/ihaeyong/Soft-TF.git and (2) CIL based on language models  [^2] https://github.com/ihaeyong/LLM-Soft-TF.git.
-5.1 Experimental Settings
-Datasets.
-(1) CL based on ViT models: we evaluate our method mainly on 1) 10/20-Split-CIFAR100 (Krizhevsky et al., 2009), constructed by splitting the 100 classes into 10 tasks/20 tasks.
-2) 10-Split-TinyImageNet (Abai and Rajmalwar, 2019), constructed by splitting the 200 classes into 10 tasks.
-3) 10-Split-ImageNet-R (Hendrycks et al., 2021), constructed by splitting the 200 classes into 10 tasks.
-To show our effectiveness, we additionally compare our method with the baselines on 5-Split-CUB200 and 10-Split-TinyImageNet.
-(2) CL based on language models: we test our method on the widely adopted continual learning benchmarks for language models following de Masson D’Autume et al.
-(2019), which use five text classification datasets (Chen et al., 2020; Zhang et al., 2015) including AG News (news classification), Yelp (sentiment analysis), DBPedia (Wikipedia article classification), Amazon (sentiment analysis) and Yahoo Answers (Q&A classification).
-Baselines.
-To validate the powerfulness of our method in (1) CL based on ViT models, we compare our results with various CIL baselines including, L2P (Wang et al., 2022c), DualPrompt (Wang et al., 2022b), and DualPrompt-PGP (Qiao et al., 2024).
-Moreover, to show the effectiveness of our method in (2) CL based on language models, we set MBPA++ (de Masson D’Autume et al., 2019), IDBR (Huang et al., 2021), SLM (Bohao et al., 2024), and DualPrompt (Wang et al., 2022b) as baselines.
-Table 2: Performances of Language Class Incremental Learning (CIL, Bert) in terms of Accuracy, Forgetting, Trainable parameters, and Time complexity on Bert benchmark datasets.
-“DR”: data replay.
-$\ast$ denotes our reproduced results.
-Order: 4 Order: 5 Order: 6 Order: 7 Average Bert-base (110.0M)
-Method DR Task ID ACC/Forget ACC/Forget ACC/Forget ACC/Forget ACC/Forget #\#Tr.Params Tr./Test[sec]
-MBPA++ $\checkmark$ L.Adapter 74.9 / 3.71 73.1 / 2.76 74.9 / 2.89 74.1 / 3.00 74.3 / 3.09 110.0M 30.12K / 87.43
-IDBR $\checkmark$ L.Adapter 75.9 / 3.46 76.2 / 2.33 76.4 / 2.88 76.7 / 2.90 76.3 / 2.89 110.0M 27.73K / 86.22
-SLM Retrieval 78.2 / 0.70 77.8 / 0.70 78.0 / 0.70 78.2 / 0.70 78.1 / 0.70 7.32M 17.51K / 73.36
-DualPrompt-L[3,4,5] Prompt ID 69.7 / 0.02 67.5 / 0.04 69.1 / 0.04 68.8 / 0.03 68.9 / 0.03 0.24M 13.86K / 70.14
-DualPrompt-L[10,11,12] Prompt ID 67.2 / 0.06 65.1 / 0.09 66.5 / 0.08 66.3 / 0.07 66.3 / 0.08 0.24M 13.86K / 70.14
-+ Adapter-L[all], r=16 Prompt ID 75.0 / 0.00 75.1 / 0.00 75.1 / 0.00 75.1 / 0.00 75.1 / 0.00 7.54M 19.45K / 75.34
-+ LoRA-L[all], r=96 Prompt ID 78.9 / 0.00 78.9 / 0.00 78.9 / 0.00 78.9 / 0.00 78.9 / 0.00 7.28M 21.57K / 76.61
-+ Soft-TF-L[3,4,5] (SOTA) Prompt ID 79.3 / 0.00 79.6 / 0.00 79.4 / 0.00 79.4 / 0.00 79.4 / 0.00 7.23M 15.51K / 71.12
-+ Soft-TF-L[10,11,12] Prompt ID 79.5 / 0.00 79.3 / 0.00 79.3 / 0.00 79.5 / 0.00 79.4 / 0.00 7.23M 15.53K / 71.12
-Task Inference.
-At the inference time, we infer task identity for arbitrary pieces of task samples $\bm{x}$ for finding the proper task nuances and demonstrating full fine-tuning results.
-We summarize the following two methods:
-  • Prompt ID: For a test example $\bm{x}$, we simply choose the best matched task index via $\text{argmin}_{t}\gamma(q(\bm{x}),\bm{k}_{t})$.
-[…]
-Table 3: Performances of Vision Class Incremental Learning (CIL, ViTs) in terms of Pretained-dataset (ImageNet-21K, SAM, DINO) and Task-IDs on 10-Split-CIFAR100 and 5-Split-CUB200.
-10-Split-CIFAR100 5-Split-CUB200
-Method Pretrained-dataset Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
-DualPrompt ImageNet-21K Prompt ID 86.50 5.77 82.02 4.23
-DualPrompt-PGP ImageNet-21K Prompt ID 86.92 5.35 82.46 3.76
-DualPrompt SAM Prompt ID 86.11 6.08 82.02 4.73
-DualPrompt DINO Prompt ID 64.18 23.81 50.88 10.10
- +Soft-TF ImageNet-21K Prompt ID 92.42 2.44 76.17 9.04
- +Soft-TF (SOTA) ImageNet-21K Gradient ID 97.87 0.21 87.93 0.66
- +Soft-TF (SOTA) SAM Gradient ID 97.87 0.21 87.93 0.66
- +Soft-TF DINO Gradient ID 84.50 12.27 69.79 10.93
-Upper-Bound of Soft-TF - - 93.90 - 85.56 -
-Refer to caption Refer to caption
-(a) L2P-PGP v.s. Soft-Transformer (b) DualPrompt v.s. Soft-Transformer
-Figure 5: Layer-wise(L[$\ast$]) Performances of Soft-TF on 10-Split-CIFAR100. Note that L[9,10,11] denotes Soft-TransFormer of 9, 10, 11 Layers.
-5.2 Performances
-Performances of Soft-TF on Vision CIL.
-We compare our Soft-TransFormers (Soft-TF) with state-of-the-art CIL baselines, as shown in Table 1.
-Our Soft-TF significantly outperforms all baselines: Adapter and LoRA and upper-bounds of Soft-TF, including L2P and DualPrompt, in accuracy and forgetting measurements.
-The time complexity of training and testing does not seem to be a significant issue under the same trainable parameters.
-The performance gain of Soft-TF is especially notable in DualPrompt-based learning compared to L2P, suggesting the importance of global prompt-tuning and multi-head attention prompt-tuning in DualPrompt.
-Additionally, task-identity inference using Gradient-ID is crucial for achieving full fine-tuning results in CIL.
-Performances of Soft-TF on Language CIL.
-Our Soft-TF significantly outperforms all baselines including MBPA++, IDBR, SLM, and DualPrompt in accuracy and forgetting measurements, as shown in Table 2.
-Replay-based methods, i.e., MBPA++ and IDBR, carry a computational burden since they replay the previous task’s samples and fine-tune 110.0M parameters.
-SLM shows inefficient time complexity since it takes two stages (warmup and fine-tuning) in the training step and the retrieval process in the inference step.
-Moreover, Soft-TF performs better than LLM-based fine-tuning baselines such as Adapter and LoRA under the same trainable parameter settings.
-Soft-TF trains and tests faster than LLM-based baselines since Soft-TF fine-tunes sparse layers (i.e., [3,4,5] layers).
-Note that LoRA failed to fine-tune onto sparse [3,4,5] layers.
-5.3 Ablation Studies
-Well-initialized LTH (WLTH) on Vision CIL.
-To demonstrate the efficacy of our proposed method on Well-initialized Lottery Ticket Hypothesis (WLTH) backbones, we evaluate our Soft-TransFormers (Soft-TF) by extending two distinct pre-trained models, ViT-DINO and ViT-SAM (Caron et al., 2021; Chen et al., 2021).
-As shown in Table 3, we tested our method on the 10-Split-CIFAR100 and 5-Split-CUB200 datasets using three pre-trained ViTs: ImageNet-21K, DINO, and SAM, further validating the effectiveness of our approach on non-ImageNet datasets (Krizhevsky et al., 2009; Wah et al., 2011).
-Surprisingly, when initialized with ImageNet-21K and SAM, DualPrompt-Soft-TF with ImageNet-21K achieved the same performance levels.
-Moreover, DualPrompt-Soft-TF outperformed all baselines, i.e., DualPrompt-PGP, on both benchmark datasets, indicating that well-initialized weights provide better generalization in continual learning scenarios.
-Random initialization.
-Random initialization of Soft-Transformer’s weights plays a critical role when leveraging well-pretrained models like Vision Transformers (ViTs).
-The optimal training point is the parameters of a well-pretrained model.
-Among the initialization methods, Uniform initialization for Soft-TransFormer satisfies this requirement effectively.
-To validate these claims, we analyze the impact of common random initialization methods, including Xavier, Kaiming, Normal, and Uniform Initialization, as shown in Table 4.
-The results demonstrate that the same well-initialization point leads to independent optimal task performance, particularly with Gradient ID inference.
-Furthermore, this ablation study strengthens our Soft-TF with state-of-the-art-perfomances inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH).
-Table 4: Random initialized Performances of Class Incremental Learning (CIL, ViTs) in terms of accuracy and forgetting on 10-Split-CIFAR100.
-Note ”w/o FF” denotes ”Soft fine-tuning without FeedForward (FF)” networks.
-DualPrompt’s parameters are pretrained on ImageNet-21K.
-10-Split-CIFAR100
-DualPrompt Task ID Random Init. ACC($\uparrow$) Forget($\downarrow$)
- +Soft-TF-L[10,11,12] w/o FF Prompt ID Xavier 90.59 3.85
- +Soft-TF-L[10,11,12] w/o FF Prompt ID Kaiming 90.72 3.63
- +Soft-TF-L[10,11,12] w/o FF Prompt ID Normal 90.45 3.78
- +Soft-TF-L[10,11,12] w/o FF Prompt ID Uniform(1.0, 1.0) 92.35 2.98
- +Soft-TF-L[10,11,12] w/o FF Gradient ID Uniform(1.0, 1.0) 98.05 0.25
-Upper-Bound of Soft-TF 93.90 -
-Layer-wise Inspections.
-We analyze the layer-wise performance of Soft-Transformer with respect to L2P and DualPrompt on the 10-Split-CIFAR100 dataset to identify the optimal configurations, as shown in Figure 5.
-Our observations reveal that the global prompt in DualPrompt influences Soft-Transformer’s performance differently in L2P and DualPrompt settings.
-In L2P-PGP, the best performance was achieved with Soft-TransFormers applied to the lower layers ((a) L2P-Soft-TF-L[1,2]-PGP), whereas in DualPrompt, the higher layers ((b) DualPrompt-Soft-TF-L[10,11,12]) yielded the best results.
-Notably, DualPrompt-Soft-TF-L[10,11,12] without PGP demonstrated impressive performance, achieving almost zero forgetting (0.21).
-These findings suggest that our approach could significantly enhance the effectiveness of large-scale Transformer models in continual learning scenarios.
 Refer to caption Refer to caption
@@
 Attention Layer.
-We compare attention maps from the 12th attention layer (L[12]) of two different models, DualPrompt and Soft-TF, applied to category 7 of the 10-Split-CIFAR100 dataset, as shown in Figure 6.
+We compare attention maps from the 12th attention layer (L[12]) of two different models, DualPrompt and Soft-TF, applied to category 7 of the 10-Split-CIFAR100 dataset, as shown in Figure˜7.
 While DualPrompt distributes its attention more uniformly, Soft-TF prioritizes specific patches of the input image, likely those more relevant to Category 7.
@@
 6 Conclusion
-Inspired by Well-initialized Lottery Ticket Hypothesis (WLTH), which provides suboptimal fine-tuning solutions, we proposed a novel fully fine-tuned continual learning (CL) method referred to as Soft-TransFormers (Soft-TF).
-Soft-TF sequentially learns and selects an optimal soft-network for each task.
-During sequential training in CL, an well-initialized Soft-TF mask optimizes the weights of sparse layers to obtain task-adaptive soft (real-valued) networks, while keeping the well-pre-trained layer parameters frozen.
-In inference, the identified task-adaptive network of Soft-TF masks the parameters of the pre-trained network, mapping to an optimal solution for each task and minimizing Catastrophic Forgetting (CF) - the soft-masking preserves the knowledge of the pre-trained network.
-Extensive experiments on the Vision Transformer (ViT) and the Language Transformer (Bert) demonstrated the effectiveness of Soft-TF, achieving state-of-the-art performance across Vision and Language CIL scenarios.
+Inspired by the Well-initialized Lottery Ticket Hypothesis (WLTH), we introduce Soft-Transformers (Soft-TF), a parameter-efficient framework for continual learning that leverages soft, real-valued subnetworks over a frozen pre-trained Transformer.
+Instead of relying on manually designed prompts or adapters, Soft-TF learns task-specific multiplicative masks applied to the key, query, value, and output projections in self-attention.
+These masks enable smooth and stable task adaptation while preserving shared representations.
+Combined with a lightweight dual-prompt mechanism, Soft-TF maintains strong knowledge retention and mitigates Catastrophic Forgetting (CF).
+Across multiple continual learning benchmarks, Soft-TF achieves state-of-the-art performance, consistently outperforming prompt-based, adapter-based, and LoRA-style baselines while requiring minimal additional parameters.
 Appendix A Appendix
@@
 To bound the right-hand side we rely on the following lemma:
-Lemma 1.
+Lemma A.1.
 Let $\bm{v}_{1},\cdots,\bm{v}_{T}$ be an arbitrary sequence of vectors. Any algorithm with an well initialization (pre-trained model) $\bm{w}^{(1)}\neq\bm{0}$ and an update rule of the form
@@
 where $\bm{w}_{m}\neq\bm{w}_{q}$ since $\bm{m}$ is learnable parameters in Soft-TransFormers.
-Specifically, we could assume that $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$ and $\bm{w}_{p}=(\bm{w}^{Q}\odot\bm{1}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{1}^{K})^{T}$ of Equation 6.
-Theorem 2.
+Specifically, we could assume that $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$ and $\bm{w}_{p}=(\bm{w}^{Q}\odot\bm{1}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{1}^{K})^{T}$ of LABEL:eq:lr_params.
+Theorem A.2.
 […]
 $\frac{1}{T}\sum_{t=1}^{T}\left<\bm{w}^{(t)}-\bm{w}^{\ast},\bm{v}_{t}\right>\leq\frac{B_{m}\rho}{\sqrt{T}}<\frac{B_{p}\rho}{\sqrt{T}}<\frac{B\rho}{\sqrt{T}}.$
-Proof  Using algebraic manipulations (completing the square), we obtain:
+Refer to caption Refer to caption
+(a) L2P-PGP v.s. L2P+Soft-TF (b) DualPrompt v.s. DualPrompt+Soft-TF
+Figure 8: Layer-wise(L[$\ast$]) Performances of Soft-TF on 10-Split-CIFAR100. Note that L[10,11,12] denotes Soft-TransFormer of 10, 11, 12 Layers.
+Proof.
+Using algebraic manipulations (completing the square), we obtain:
 […]
@@
 The second part follows by upper bounding $\bm{w}||$ by $B$, $\bm{v}_{t}||$ by $\rho$, deciding by $T$, and plugging in the value of $\eta$.
-In terms of Soft-TransFormers $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$ of Equation 6, we have
+Table 5: Performances of Class Incremental Learning (CIL) in terms of accuracy and forgetting on $10/20$-Split-CIFAR100 and 10-Split-ImageNet-R.
+Exemplar means the total buffer size for rehearsal methods.
+ViT-B/16 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
+Method #Tr.Params. Task ID ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec]
+BiC 5,000 - 81.42 17.31 73.02 6.23 64.63 22.25
+DER++ 5,000 - 83.94 14.55 - - 66.73 20.67
+iCaRL 5,000 - 66.00 5.33 78.02 5.80 - -
+DER+MCG 2,000 - 67.62 14.64 65.84 13.72 - -
+BiC 1,000 - 66.11 35.24 63.12 21.89 52.14 36.70
+DER++ 1,000 - 61.06 39.87 - - 55.47 34.64
+iCaRL 1,000 - 61.25 14.19 71.32 15.98 - -
+FT - - 33.61 86.87 33.52 53.69 28.87 63.80
+EWC - - 47.01 33.27 36.73 35.19 35.00 56.16
+LWF - - 60.69 27.77 39.12 57.91 38.54 52.37
+DualPrompt 0.03M Prompt ID 86.50 / 5.77 12.12K / 76 82.98 / 8.20 11.60K / 78 68.13 / 4.46 13.10K / 47
++ PGP 0.03M Prompt ID 86.92 / 5.35 12.21K / 76 83.74 / 7.91 13.12K / 78 63.34 / 4.53 13.33K / 47
++ Soft-TF-L[3,4,5] 6.93M Prompt ID 91.77 / 3.37 13.87K / 80 94.43 / 2.02 15.60K / 104 74.70 / 6.46 15.35K / 52
++ Soft-TF-L[3,4,5] 6.93M Gradient ID 93.76 / 1.83 13.87K / 130 95.38 / 1.73 15.60K / 163 82.15 / 2.20 15.35K / 82
++ Soft-TF-L[10,11,12] 6.93M Prompt ID 92.35 / 2.98 13.87K / 80 97.40 / 0.57 15.60K / 104 76.62 / 5.30 15.35K / 52
++ Soft-TF-L[10,11,12] 6.93M Gradient ID 97.87 / 0.21 13.87K / 130 99.05 / 0.24 15.60K / 163 82.38 / 0.59 15.35K / 82
+Upper-Bound of DualPrompt 0.03M - 90.85 - 90.85 - 79.13 -
+Upper-Bound of Soft-TF 6.93M - 95.90 - 95.90 - 80.21 -
+In terms of Soft-TransFormers $\bm{w}_{m}=(\bm{w}^{Q}\odot\bm{m}^{Q})\cdot\bm{x}\bm{p}^{T}\cdot(\bm{w}^{K}\odot\bm{m}^{K})^{T}$, we have
 […]
@@
 $\frac{1}{T}\sum_{t=1}^{T}\left<\bm{w}^{(t)}-\bm{w}^{\ast},\bm{v}_{t}\right>\leq\frac{B_{m}\rho}{\sqrt{T}}<\frac{B_{p}\rho}{\sqrt{T}}<\frac{B\rho}{\sqrt{T}}.$
-Refer to caption Refer to caption
-(a) DualPrompt-PGP v.s. Soft-TF (b) DualPrompt-WSN v.s. Soft-TF
-[…]
 A.2 Experimental Details
-For fair comparisons with the baselines (Wang et al., 2022c, b; Qiao et al., 2024), we use ViT B/16 (Dosovitskiy et al., 2020) pre-trained on ImageNet-21K as our image encoder, which is kept frozen during training.
+For fair comparisons with the baselines (Wang et al., 2022d, c; Qiao et al., 2024), we use ViT B/16 (Dosovitskiy et al., 2021) pre-trained on ImageNet-21K as our image encoder, which is kept frozen during training.
 We train and test on a single Quadro RTX 8000-48GB GPU for baselines and our Soft-TransFormers with Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
-We adhere to the experimental settings outlined by Qiao et al.
-(2024) to validate our method’s effectiveness.
+We adhere to the experimental settings outlined by (Qiao et al., 2024) to validate our method’s effectiveness.
 When comparing our approach with L2P-PGP and Soft-Transformer on the 10/20-Split-CIFAR100 and 10-Split-TinyImageNet datasets, we train the network for 5 epochs with a batch size of 16 and set the prompt length to 5.
@@
 Additionally, in all benchmark data sets, the general prompt length is set to 5, and the location inserted into the prompt is kept consistent.
-CLIP on CIL and TIL.
-We conduct our experiments on the 10-Split-CIFAR100 dataset under both Class Incremental Learning (CIL) and Task Incremental Learning (TIL) settings, as shown in Table 5.
-The results demonstrate that CLIP-Prompt-Soft-TF-L[1-12] significantly improves performance in both settings, indicating that our Soft-TF with Gradient ID is also effective in vision-language models, thereby broadening its applicability.
-Table 5: Comparisons of Soft-TF with baselines based on CLIP model on 10-Split-CIFAR100. $\ast$ denotes our reproduced results.
-Class Incremental Task Incremental
-Method Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
-CLIP Prompt ID 73.76 5.60 92.69 2.34
-CLIP-PGP Prompt ID 79.47 4.23 93.00 1.58
-CLIP Prompt ID 74.60 7.75 93.59 2.80
-CLIP-PGP Prompt ID 74.63 7.76 93.67 2.83
-CLIP-Prompt Prompt ID 70.27 12.95 93.36 3.07
-CLIP-Prompt-Soft-TF-L[3,4,5] Prompt ID 71.58 7.73 95.29 1.12
-CLIP-Prompt-Soft-TF-L[3,4,5] Gradient ID 76.77 5.59 95.29 1.12
-CLIP-Prompt-Soft-TF-L[1-12] Prompt ID 72.28 3.44 96.83 0.44
-CLIP-Prompt-Soft-TF-L[1-12] (SOTA) Gradient ID 85.90 3.07 96.83 0.44
-For CLIP-PGP and Soft-TransFormers, we configure a single trainable image prompt that is shared across all tasks within the vision encoder.
-For the text encoder, following the approach of Qiao et al.
-(2024), we set a trainable text prompt for each class, which is only trained on the corresponding task.
-In our comparisons with CLIP-PGP and Soft-TransFormers on the 10-Split-CIFAR100 dataset, we set the image prompt length to 5, the number of epochs to 5, and the batch size to 32.
 Table 6: Performances of Class Incremental Learning (CIL) in terms of accuracy and forgetting on $10/20$-Split-CIFAR100 and 10-Split-ImageNet-R.
-Exemplar means the total buffer size for rehearsal methods.
-10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
-Method Exemplar Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
-BiC 5,000 - 81.42 17.31 73.02 6.23 64.63 22.25
-DER++ 5,000 - 83.94 14.55 - - 66.73 20.67
-iCaRL 5,000 - 66.00 5.33 78.02 5.80 - -
-DER+MCG 2,000 - 67.62 14.64 65.84 13.72 - -
-BiC 1,000 - 66.11 35.24 63.12 21.89 52.14 36.70
-DER++ 1,000 - 61.06 39.87 - - 55.47 34.64
-iCaRL 1,000 - 61.25 14.19 71.32 15.98 - -
-FT - - 33.61 86.87 33.52 53.69 28.87 63.80
-EWC - - 47.01 33.27 36.73 35.19 35.00 56.16
-LWF - - 60.69 27.77 39.12 57.91 38.54 52.37
-L2P - Prompt ID 83.77 6.63 71.29 13.96 60.44 9.00
-L2P-PGP - Prompt ID 84.34 5.59 76.12 13.26 61.40 8.03
-L2P-PGP-Soft-TF - Prompt ID 86.26 4.79 76.17 15.77 69.80 5.13
-L2P-PGP-Soft-TF - Gradient ID 86.46 4.87 77.67 15.84 69.56 5.28
-DualPrompt - Prompt ID 86.50 5.77 82.98 8.20 68.13 4.46
-DualPrompt-Soft-TF-L[3,4,5] - Prompt ID 91.77 3.37 94.43 2.02 74.70 6.46
-DualPrompt-Soft-TF-L[3,4,5] - Gradient ID 93.76 1.83 95.38 1.73 82.15 2.20
-DualPrompt-WSN-L[10,11,12], c=80.0 $\%$ - Gradient ID 97.41 0.18 90.25 9.08 74.83 0.91
-DualPrompt-WSN-L[10,11,12], c=81.0 $\%$ - Gradient ID 97.50 0.21 96.72 2.21 74.21 1.41
-DualPrompt-WSN-L[10,11,12], c=82.0 $\%$ - Gradient ID 97.67 0.27 96.44 1.62 75.02 0.92
-DualPrompt-WSN-L[10,11,12], c=87.0 $\%$ - Gradient ID 97.51 0.27 97.68 0.75 77.96 1.02
-DualPrompt-WSN-L[10,11,12], c=90.0 $\%$ - Gradient ID 97.46 0.38 98.09 0.65 78.80 0.47
-DualPrompt-Soft-TF-L[10,11,12] - Gradient ID 97.87 0.21 99.05 0.24 82.38 0.59
-DualPrompt-PGP - Prompt ID 86.92 5.35 83.74 7.91 69.34 4.53
-DualPrompt-PGP-Soft-TF-L[3,4,5] - Prompt ID 92.41 2.44 95.14 1.90 74.65 4.39
-DualPrompt-PGP-Soft-TF-L[3,4,5] - Gradient ID 92.92 2.34 95.89 1.64 81.45 2.89
-Upper-Bound of DualPrompt - - 90.85 - 90.85 - 79.13 -
-Upper-Bound of Soft-TF - - 93.90 - 93.90 - 80.21 -
-Table 7: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, training, and test time on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R.
-Note ”w/o FF” denotes ”Soft finetuning without FeedForward (FF)” networks.
-Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
-DualPrompt # Train Params. Task ID Train (sec.) Test (sec.) Train (sec.) Test (sec.) Train (sec.) Test (sec.)
-DualPrompt 0.03M Prompt ID 12.12K 76 11.60K 78 13.10K 47
-  PGP 0.03M Prompt ID 12.21K 76 13.12K 78 13.33K 47
-  Soft-TF-L[12] w/ only ATTN 1.76M Gradient ID 12.18K 129 13.30K 113 13.35K 65
-  Soft-TF-L[12] w/ only ATTN 1.76M Prompt ID 12.18K 78 13.30K 80 13.35K 48
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID 12.24K 103 13.40K 132 13.42K 66
-  Soft-TF-L[11,12] w/o FF 4.62M Gradient ID 12.95K 115 14.38K 146 14.23K 73
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID 13.71K 130 15.51K 163 15.08K 82
-  Soft-TF-L[10,11,12] w/o FF 6.93M Prompt ID 13.87K 80 15.60K 104 15.35K 52
-  LoRA-L[10,11,12] w/o FF, r=4 0.06M Prompt ID 11.95K 77 11.71K 79 14.34K 48
-  LoRA-L[10,11,12] w/o FF, r=24 0.32M Prompt ID 12.03K 78 15.10K 100 15.02K 50
-  LoRA-L[10,11,12] w/o FF, r=500 6.93M Prompt ID 13.24K 79 15.89K 105 15.09K 53
-  Adapter-L[10,11,12] w/ FF, r=1 0.09M Prompt ID 12.44K 84 12.40K 81 14.40K 50
-  Adapter-L[10,11,12] w/ FF, r=4 0.36M Prompt ID 12.80K 85 15.35K 105 14.68K 51
-  Adapter-L[10,11,12] w/ FF, r=300 6.93M Prompt ID 13.66K 88 15.72K 106 15.50K 53
-Table 8: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, Accuracy, and Forget on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R.
-Note ”w/o FF” denotes ”Soft finetuning without FeedForward (FF)” networks.
-Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
-DualPrompt # Train Params. Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
-DualPrompt 0.03M Prompt ID 86.50 5.77 82.98 8.20 68.13 4.46
-  PGP 0.03M Prompt ID 86.92 5.35 83.74 7.91 69.34 4.53
-  Soft-TF-L[12] w/ only ATTN 1.76M Gradient ID 97.17 0.40 98.09 0.54 72.31 3.94
-  Soft-TF-L[12] w/ only ATTN 1.76M Prompt ID 94.59 1.12 96.96 1.02 71.13 4.93
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID 96.84 0.55 97.81 0.57 81.18 1.31
-  Soft-TF-L[11,12] w/o FF 4.62M Gradient ID 97.58 0.34 98.65 0.43 83.09 0.42
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID 98.05 0.25 98.96 0.23 83.70 0.53
-  Soft-TF-L[10,11,12] w/o FF 6.93M Prompt ID 92.35 2.98 97.40 0.57 76.62 5.30
-  LoRA-L[10,11,12] w/o FF, r=4 0.06M Prompt ID 82.19 4.33 93.74 2.07 70.91   9.11
-  LoRA-L[10,11,12] w/o FF, r=24 0.32M Prompt ID 86.77 4.27 95.65 1.04 69.81 10.30
-  LoRA-L[10,11,12] w/o FF, r=500 6.93M Prompt ID 82.00 4.33 92.14 2.02 43.51 13.21
-  Adapter-L[10,11,12] w/ FF, r=1 0.09M Prompt ID 86.38 4.87 85.61 5.04 70.95 4.31
-  Adapter-L[10,11,12] w/ FF, r=4 0.36M Prompt ID 86.53 4.52 85.66 5.00 70.82 4.90
-  Adapter-L[10,11,12] w/ FF, r=300 6.93M Prompt ID 86.45 4.61 84.75 5.11 70.55 4.74
-Training & Test Time.
-To clearly illustrate the time complexity of Soft-TF, we present the training and testing times for 10/20-Split-CIFAR100 and 10-Split-ImageNet-R, as shown in Table 7.
-As the number of trainable parameters in Soft-TF increases, training and testing time complexities grow accordingly.
-While the testing time complexity of Gradient ID increased by approximately 1.6 times across the three benchmark datasets, it consistently improved task performance on all benchmarks.
-The corresponding performance metrics are detailed in Table 8.
-We also investigate the most parameter-efficient and gradient-based task inference methods, as shown in Table 9 and Table 10.
-Our findings reveal that the 3-shot Gradient ID inference cost (using samples within a mini-batch) with the last layer (Soft-TF-L[12]) is approximately 1.1 times that of Prompt ID, maintaining comparable efficiency while delivering superior performance.
-Note that m-batch denotes mini-batch.
-Table 9: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, training, and test time on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R.
-Note ”w/o FF” denotes ”Soft finetuning without FeedForward (FF)” networks.
-Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
-DualPrompt # Train Params. Task ID Train (sec.) Test (sec.) Train (sec.) Test (sec.) Train (sec.) Test (sec.)
-DualPrompt 0.03M Prompt ID 12.12K 76 11.60K 78 13.10K 47
-  PGP 0.03M Prompt ID 12.21K 76 13.12K 78 13.33K 47
-  Soft-TF-L[12] w/o FF 2.31M Prompt ID 12.24K 79 13.40K 80 13.42K 48
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 3-shot 12.24K 88 13.40K 90 13.42K 57
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 5-shot 12.24K 94 13.40K 98 13.42K 61
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 7-shot 12.24K 95 13.40K 108 13.42K 62
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID, m-batch 12.24K 103 13.40K 132 13.42K 66
-  Soft-TF-L[10,11,12] w/o FF 6.93M Prompt ID 13.87K 80 15.60K 104 15.35K 52
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 3-shot 13.71K 96 15.51K 106 15.08K 63
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 5-shot 13.71K 96 15.51K 109 15.08K 74
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 7-shot 13.71K 106 15.51K 119 15.08K 75
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, batch 13.71K 130 15.51K 163 15.08K 82
-Table 10: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, Accuracy, and Forget on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R.
-Note ”w/o FF” denotes ”Soft finetuning without FeedForward (FF)” networks.
-Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
-DualPrompt # Train Params. Task ID ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$) ACC($\uparrow$) Forget($\downarrow$)
-DualPrompt 0.03M Prompt ID 86.50 5.77 82.98 8.20 68.13 4.46
-  PGP 0.03M Prompt ID 86.92 5.35 83.74 7.91 69.34 4.53
-  Soft-TF-L[12] w/o FF 2.31M Prompt ID 91.83 2.99 96.43 1.00 72.45 5.32
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 3-shot 93.12 1.82 96.43 1.00 73.55 4.80
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 5-shot 96.13 0.58 96.43 1.00 75.04 4.49
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID, 7-shot 96.51 0.65 96.43 1.00 76.34 4.75
-  Soft-TF-L[12] w/o FF 2.31M Gradient ID, batch 96.84 0.55 97.81 0.57 81.18 1.31
-  Soft-TF-L[10,11,12] w/o FF 6.93M Prompt ID 92.35 2.98 97.40 0.57 74.62 5.30
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 3-shot 93.92 1.62 97.40 0.57 74.99 4.21
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 5-shot 97.37 0.53 97.40 0.57 77.40 2.91
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, 7-shot 97.76 0.51 97.40 0.57 79.33 3.57
-  Soft-TF-L[10,11,12] w/o FF 6.93M Gradient ID, m-batch 98.05 0.25 98.96 0.23 83.70 0.53
-Soft & Subnetworks have been explored in continual learning through the following notable approaches such as supermasks (Wortsman et al., 2020), WSN (Kang et al., 2022), SoftNet (Kang et al., 2023), and FSO (Kang et al., 2024a).
-All prior works have primarily demonstrated the effectiveness of sparse solutions in fully connected networks and Convolutional Neural Networks (CNNs).
-Sparse fine-tuning mechanisms have not been investigated for pre-trained transformers, such as ViT, and remain unexplored.
-To discover the competitive sparseness in Transformers, we detail the WSN-style task-adaptive fine-tuning and the learnable soft-networks $\bm{m}$ of Transformers, presenting these adaptations for the first time with empirical observations.
-Here, the Soft-TF originates from learned parameters distributed with $\mu\approx 1.0$ & various variances, as stated in Figure 3: we observed that each task solution has its unique variances which help infer task-ID in gradient step.
-To demonstrate the effectiveness of Soft-TF, Figure 7(a) presents a radar chart comparing DualPrompt-PGP and Soft-TransFormers across four benchmark datasets with various sparseness (Figure 7(b) & Table 6.): we compare the performance of DualPrompt-WSN and Soft-TF using accuracy across tasks on the 10-Split CIFAR100 dataset.
-The performances of Soft-TF provide the upper bound of WSN, as shown in Figure 8(a).
-In Vision Transformers, Figure 8(b) demonstrates the exsistances of a smooth solution of SoftNet.
-Comparisions of Soft-TF with LLMs.
-To demonstrate the effectiveness of Soft-TF, we compare Soft-TF against LLM fine-tuning methods such as Adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021), as shown in Table 8.
-Under identical experimental conditions—including trainable model parameters (6.9M per task), layers (L[10,11,12]), and Prompt ID—Soft-TF outperformed other LLM-based fine-tuning approaches.
-The results highlight that directly updating well-pretrained model parameters and prompt-tuning via Soft-TF is more effective than combining representations through LoRA or learning representations with Adapters.
-Furthermore, we observed that Soft-TF and the other methods exhibited comparable training and testing time complexity for the same number of trainable parameters.
-Notably, single-layer fine-tuning using Soft-TF (with L[12]) surpasses the performance of the baselines.
-These findings firmly establish Soft-TF as the most competitive approach among strong LLM baselines (Adapters and LoRA) in the continual learning (CIL) scenario.
-Sparsity of Transformer.
-We inspect the sparse solution through WSN as shown in Table 6, Table 11, and Table 12.
-We found a suboptimal sparse solution (c=87.0 % on 10-Split-TinyImageNet) with minimal CF through the inspections, as stated in Table 11.
-This demonstrates the Lottery Ticket Hypothesis (LTH) in transformers, a competitive sparse subnetwork in DenseNetwork.
-In addition, DualPrompt is the lower-bound while DualPrompt-Soft-TF-$\ast$ is the upper-bound, close to the optimal performances.
-Refer to caption Refer to caption
-(a) WSN-TransFormers v.s. Soft-TransFormers (b) WSN v.s. SoftNet
-Figure 8: Comparisions of Soft-TransFormers with Subnetworks on 10-Split-CIFAR100. Note that L[10,11,12] denotes the fine-tuning layers of 10, 11, and 12.
-Table 11: Performances of Subnetworks (WSN) in Class Incremental Learning (CIL) on 10-Split-TinyImageNet.
-TinyImageNet
-Method Pretrained-Dataset Task ID ACC($\uparrow$) Forget($\downarrow$)
-DualPrompt - Prompt ID 86.50 5.77
-DualPrompt-WSN-L[10,11,12] C=80.0% Gradient ID 89.95 0.98
-DualPrompt-WSN-L[10,11,12] C=81.0% Gradient ID 89.99 1.06
-DualPrompt-WSN-L[10,11,12] C=82.0% Gradient ID 89.59 1.18
-DualPrompt-WSN-L[10,11,12] C=83.0% Gradient ID 90.60 0.72
-DualPrompt-WSN-L[10,11,12] C=85.0% Gradient ID 90.09 1.07
-DualPrompt-WSN-L[10,11,12] C=87.0% Gradient ID 91.91 0.38
-DualPrompt-WSN-L[10,11,12] C=90.0% Gradient ID 91.28 0.42
-DualPrompt-WSN-L[10,11,12] C=93.0% Gradient ID 91.41 0.40
-DualPrompt-WSN-L[10,11,12] C=95.0% Gradient ID 90.91 0.83
-DualPrompt-Soft-TF-L[10,11,12] - Gradient ID 97.87 0.21
-Table 12: Performances of Class Incremental Learning (CIL) in terms of Soft parameters, Accuracy, and Forget on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R.
-Note ”w/ only ATTN” denotes ”Soft finetuning only with attention (ATTN)” and FeedForward ($F$) networks.
-Method ViT-B/12 (85.8M) 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
-DualPrompt #Tr. Params. Task ID ACC Forget ACC Forget ACC Forget
-DualPrompt 0.03M Prompt ID 86.50 5.77 82.98 8.20 68.13 4.46
-  PGP 0.03M Prompt ID 86.92 5.35 83.74 7.91 69.34 4.53
-  Soft-TF-L[12] w/ only ATTN 1.76M Gradient ID 97.17 0.40 98.09 0.54 72.31 3.94
-  Soft-TF-L[12] w/ only ATTN, c=90% 1.58M Gradient ID 96.81 0.61 97.51 1.68 71.67 3.94
+ViT-B/16 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
+Method #Tr.Params. Task ID ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec] ACC/Forget Tr./Test[sec]
+L2P 0.03M Prompt ID 83.77 / 6.63 12.00K / 75 71.29 / 13.96 11.50K / 76 60.44 / 9.00 12.80K / 46
++ Soft-TF-L[3,4,5] 6.93M Prompt ID 86.26 / 4.79 12.85K / 78 76.17 / 15.77 13.98K /100 69.80 / 5.13 14.23K / 49
+DualPrompt 0.03M Prompt ID 86.50 / 5.77 12.12K / 76 82.98 / 8.20 11.60K / 78 68.13 / 4.46 13.10K / 47
++ Soft-TF-L[10,11,12] 6.93M Prompt ID 92.35 / 2.98 13.87K / 80 97.40 / 0.57 15.60K / 104 76.62 / 5.30 15.35K / 52
+S-Prompt 0.03M Prompt ID 87.57 / 3.63 12.12K / 76 84.90 / 7.05 11.60K / 78 74.25 / 4.73 13.10K / 47
+CODA-Prompt 0.03M Prompt ID 86.94 / 4.04 12.12K / 76 84.70 / 7.11 11.60K / 78 74.26 / 5.17 13.10K / 47
+HiDe-Prompt 0.03M Prompt ID 92.61 / 1.50 12.12K / 76 94.28 / 1.78 11.60K / 78 75.06 / 4.09 13.10K / 47
+NoRGa 0.03M Prompt ID 94.48 / 1.44 12.12K / 76 97.39 / 0.58 11.60K / 78 75.40 / 4.59 13.10K / 47
++ Soft-TF-L[10,11,12] 6.93M Prompt ID 98.25 / 0.31 13.87K / 80 98.30 / 0.30 15.60K / 104 91.94 / 0.34 15.35K / 52
+Comparisons with Prompt-based Baselines.
+We evaluate the performance of Soft-TF against representative prompt-based continual learning methods, including L2P (Wang et al., 2022d), DualPrompt (Wang et al., 2022c), CODA-Prompt (Smith et al., 2023b), S-Prompt (Wang et al., 2022a), HiDe-Prompt (Wang et al., 2023), and NoRGa (Le et al., 2024), as summarized in Table˜6.
+While these approaches demonstrate that improved prompt representations generally lead to better continual learning performance, their adaptation capacity remains fundamentally constrained by prompt-level conditioning.
+In contrast, Soft-TF consistently achieves state-of-the-art results by directly modulating attention projections through task-specific soft subnetworks, providing substantially greater representational expressiveness than prompts alone.
+These results indicate that, even when compared to strong and diverse prompt-based baselines, Soft-TF offers a more effective and scalable mechanism for continual adaptation.
+Task Inference.
+At the inference time, we infer task identity for arbitrary pieces of task samples $\bm{x}$ for finding the proper task nuances and demonstrating full fine-tuning results.
+We summarize the following two methods:
+  • Prompt ID: For a test example $\bm{x}$, we simply choose the best matched task index via $\text{argmin}_{t}\gamma(q(\bm{x}),\bm{k}_{t})$.
+[…]
+Using the Gradient ID inference method, as shown in Table˜5, Soft-TF demonstrates superior average performance and minimal forgetting compared to the Prompt ID approach, even with a slight increase in computational load at inference time.
+This method is not only valuable for applications requiring the highest level of performance but also serves as a foundational study for future research into prompt learning and task-ID inference.
+Layer-wise Inspections.
+We analyze the layer-wise performance of Soft-Transformer with respect to L2P and DualPrompt on the 10-Split-CIFAR100 dataset to identify the optimal configurations, as shown in Figure˜8.
+Our observations reveal that the global prompt in DualPrompt influences Soft-Transformer’s performance differently in L2P and DualPrompt settings.
+In L2P-PGP, the best performance was achieved with Soft-TransFormers applied to the lower layers ((a) L2P-Soft-TF-L[1,2]-PGP), whereas in DualPrompt, the higher layers ((b) DualPrompt-Soft-TF-L[10,11,12]) yielded the best results.
+Notably, DualPrompt-Soft-TF-L[10,11,12] without PGP demonstrated impressive performance, achieving almost zero forgetting (0.21).
+These findings suggest that our approach could significantly enhance the effectiveness of large-scale Transformer models in continual learning scenarios.
 Peudo Codes.
-The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm 1 and Algorithm 2.
+The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm˜1 and Algorithm˜2.
 We denote the architecture with attached prompts as $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$.
@@
 Finally, we train the two prompts, the task keys, the soft-attention parameters, and the newly-initialized classification head in an end-to-end manner.
-Algorithm 1 DualPrompt-Soft-TF at training time
-1:  Input: Pre-trained transformer-based backbone $f$, final classification layer $f_{\phi}$,
-2:       number of tasks $\mathcal{T}$, training set $\{\{\bm{x}_{i,t},y_{i,t}\}_{i=1}^{n_{t}}\}_{t=1}^{\mathcal{T}}$, G-Prompt $\bm{g}$, E-Prompt $\bm{E}=\{\bm{e}_{t}\}_{t=1}^{\mathcal{T}}$,
-3:       task keys $\bm{K}=\{\bm{k}_{t}\}_{t=1}^{T}$, soft-networks $\bm{M}=\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t}}\}_{t=1}^{\mathcal{T}}$ $start_{g},end_{g},start_{e},end_{e}$,
-4:       prompting function $f^{prompt}_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$,
-5:       number of training epochs of the $t$-th task $\mathcal{K}_{t}$.
-6:  Initialize: $\phi,\bm{g},\bm{E},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{M}},\bm{K}$
-7:  for task $t=1,\cdots,\mathcal{T}$ do
-8:  Select the task-specific E-Prompt, soft-network $\bm{e}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t}}$ and corresponding task key $\bm{k}_{t}$
-9:  Generate the prompted architecture $f_{\bm{g},\bm{e}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t}}}$: attach $\bm{g}$ and $\bm{e}_{t}$ to $start_{g}$-th to $end_{g}$-th
-10:       and $start_{e}$-th to $end_{e}$-th soft MSA layers respectively, with $f^{prompt}_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$.
-11:  for batch $\bm{e}_{s}\sim\mathcal{K}_{t}$ do
-12:   Draw a mini-batch $B=\{(\bm{x}_{i,t},y_{i,t})\}_{i=1}^{l}$
-13:   for ($\bm{x},y$) in B do
-14:    Calculate the prompted feature by
-15:    Calculate the per sample loss $\mathcal{L}_{x}$ via
-16:   end for
-17:   Update $\phi,\bm{g},\bm{E},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{M}},\bm{K}$ by back-propagation
-18:  end for
-19:  end for
-Algorithm 2 DualPrompt-Soft-TF at test time
-1:  Given components: Pre-trained transformer-based backbone $f$, trained
-2:       $\bm{K}=\{\bm{k}_{t}\}_{t=1}^{\mathcal{T}},\bm{M}=\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t}}\}_{t=1}^{\mathcal{T}},start_{g},end_{g},start_{e},end_{e},$ prompting function $f^{prompt}_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$
-3:  Input: test example $\bm{x}$ from mini-batch $\bm{b}$
-4:  Select task inference method: (1) Prompt ID or (2) Gradient ID
-5:     (1) Prompt ID:
-6:       Generate query feature $q(\bm{x})$
-7:       Matching for the index of E-Prompt via $t_{\bm{x}}=\mathrm{argmin}_{t}\gamma(q(\bm{x}),\bm{k}_{t})$
-8:     (2) Gradient ID:
-9:       Assigning each learned subnetwork $\bm{m}_{t}$ a weight $\alpha_{t}$ such that $\sum_{t}\alpha_{t}=1$ and $\alpha_{t}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1/\mathcal{T}}>0.$
-10:       Given $\bm{x}\in\bm{b}$ to classify, we can compute our loss $\mathcal{L}=\mathcal{H}(f^{prompt}_{\bm{\theta}\odot(\sum_{t}\alpha_{t}\bm{m}_{t})}(\bm{x}))$
-11:       Matching for the index of E-Prompt via $t_{\bm{x}}=\mathrm{argmin}_{t}\frac{\partial\mathcal{H}}{\partial\alpha_{t}}$
-12:  Select the task-specific E-prompt $\bm{e}_{t_{\bm{x}}}$ and learned subnetwork $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t_{\bm{x}}}$}
-13:  Generate the prompted architecture $f_{\bm{g},\bm{e}_{t_{\bm{x}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t_{\bm{x}}}}}$:
-14:       Attaching $\bm{g}$ and $\bm{e}_{t_{\bm{x}}}$ to $start_{g}$-th to $end_{g}$-th
-15:       and $start_{e}$-th to $end_{e}$-th MSA layers respectively, with $f^{prompt}_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$.
-16:  Prediction: $f_{\bm{g},\bm{e}_{t_{\bm{x}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}_{t_{\bm{x}}}}}(\bm{x})$
-Density of Parameters.
-We inspect the histogram density estimate of the last (12) layer’s parameters of DualPrompt-Soft-TF: attention of QKV ((a) ATTN.QKV) and Projection ((b) ATTN.PROJ) and multi-layer perception (MLP) of FC1 and FC2, as shown in Figure 9.
-ATTN’s QKV parameters have the largest variance among the parameter densities, while MLP-FC2’s are the smallest.
-From this observation, we conclude that fine-tuning ATTN’s QKV is required to achieve optimal task performance.
-In other words, QKV’s parameters are more critical than others.
 Refer to caption Refer to caption
@@
 Figure 9: Layer-(L[12]) Histogram Density Estimates of DualPrompt-Soft-TF’s Parameters on 10-Split-CIFAR100.
+Refer to caption Refer to caption
+(a) $\bm{\theta}^{QKV}_{Attn}$ v.s. $\bm{\theta}^{QKV}_{Attn}\odot\bm{m}^{QKV}_{Attn}$ (b) $\bm{\theta}^{O}_{Proj}$ v.s. $\bm{\theta}^{O}_{Proj}\odot\bm{m}^{O}_{Proj}$
+Refer to caption Refer to caption
+(c) $\bm{\theta}^{fc1}$ v.s. $\bm{\theta}^{fc1}\odot\bm{m}^{fc1}$ (d) $\bm{\theta}^{fc2}$ v.s $\bm{\theta}^{fc2}\odot\bm{m}^{fc2}$
+Figure 10: Layer-(L[12]) Histogram Density Estimates of Pre-trained Weight and DualPrompt-Soft-TF’s Parameters on 10-Split-CIFAR100.
+Density of Parameters.
+We inspect the histogram density estimate of the last (12) layer’s parameters of DualPrompt-Soft-TF: attention of QKV ((a) Attn.QKV) and Projection ((b) Attn.Proj) and multi-layer perception (Mlp) of fc1 and fc2, as shown in Figure˜9.
+Attn’s QKV parameters have the largest variance among the parameter densities, while MLP-FC2’s are the smallest.
+From this observation, we conclude that fine-tuning Attn’s QKV is required to achieve optimal task performance.
+In other words, QKV’s parameters are more critical than others.
 Pre-trained Parameters vs.
 Soft-TF.
-We inspect the histogram density estimate of the last (12) layer’s parameters of pre-trained model and DualPrompt-Soft-TF: attention of QKV ((a) ATTN.QKV) and Projection ((b) ATTN.PROJ) and multi-layer perception (MLP) of FC1 and FC2, as shown in Figure 10.
+We inspect the histogram density estimate of the last (12) layer’s parameters of pre-trained model and DualPrompt-Soft-TF: attention of QKV ((a) Attn.QKV) and Projection ((b) Attn.Proj) and multi-layer perception (Mlp) of fc1 and fc2, as shown in Figure˜10.
 The most parameters of DualPrompt-Soft-TF are trained around zero-values.
 Particularly, the difference between pre-trained model’s parameters and Soft-TF is distinctive at QKV module.
-Refer to caption Refer to caption
-(a) ATTN.QKV ($\bm{w}^{QKV}$ v.s. $\bm{w}^{QKV}\odot\bm{m}^{QKV}$) (b) ATTN.PROJ ($\bm{w}^{O}$ v.s. $\bm{w}^{O}\odot\bm{m}^{O}$)
-Refer to caption Refer to caption
-(c) MLP.FC1 ($\bm{w}^{FC1}$ v.s. $\bm{w}^{FC1}\odot\bm{m}^{FC1}$) (d) MLP.FC2 ($\bm{w}^{FC2}$ v.s $\bm{w}^{FC2}\odot\bm{m}^{FC2}$)
-Figure 10: Layer-(L[12]) Histogram Density Estimates of Pre-trained Weight and DualPrompt-Soft-TF’s Parameters on 10-Split-CIFAR100.
+Additional Comparisons with Prompt-based Baselines.
+As shown in Table˜7, Soft-Transformers (Soft-TF) achieve a substantial performance margin over established prompt-based continual learning methods on 10-Split-CIFAR100.
+Whereas conventional prompt learners such as L2P, DualPrompt, ESN, CODA-Prompt, and CPrompt obtain accuracies in the range of 77–87% with forgetting between 5–7%, Soft-TF reaches 92.35% accuracy and reduces forgetting to only 2.98.
+This improvement is not incremental but a clear step-change: Soft-TF reduces forgetting by nearly half while surpassing the best prompt-based accuracy by more than 5 percentage points.
+These results highlight a fundamental limitation of prompt-only mechanisms, which rely solely on input-level conditioning and thus struggle to capture fine-grained, task-specific distinctions under heterogeneous task streams.
+By contrast, Soft-TF leverages task-specific soft subnetworks that directly modulate attention projections, offering richer representational flexibility while maintaining the frozen backbone.
+The strong gains over diverse prompt baselines underscore that Soft-TF provides a more expressive and stable adaptation mechanism than prompting alone, even when evaluated under random initialization.
+Algorithm 2 DualPrompt-Soft-TF at Test Time
+1:  Given: Pre-trained $f$; trained task keys $\bm{K}=\{\bm{k}_{t}\}_{t=1}^{\mathcal{T}}$;
+2:   learned soft-networks $\bm{M}=\{\bm{m}_{t}\}_{t=1}^{\mathcal{T}}$;
+3:   prompt insertion layers $l^{start}_{g},l^{end}_{g}]$, $l^{start}_{e},l^{end}_{e}]$;
+4:   prompting function $f_{\bm{\theta}\odot\bm{m}}$.
+5:  Input: Test example $\bm{x}$ (from mini-batch $\bm{b}$)
+6:  Task inference method:
+7:   Choose either (1) Prompt ID or (2) Gradient ID
+8:  (1) Prompt ID
+9:   Compute query feature $q(\bm{x})$
+10:   Select task index:
+11:   $t_{\bm{x}}=\arg\min_{t}\gamma\big(q(\bm{x}),\bm{k}_{t}\big)$
+12:  (2) Gradient ID
+13:   Initialize task weights $\{\alpha_{t}\}_{t=1}^{\mathcal{T}}$,
+14:   with $\alpha_{t}=1/\mathcal{T}$ and $\sum_{t}\alpha_{t}=1$
+15:   Compute soft-merged network:
+16:   $\bm{m}_{\alpha}=\sum_{t=1}^{\mathcal{T}}\alpha_{t}\bm{m}_{t}$
+17:   Compute loss on $\bm{x}$:
+18:   $\mathcal{L}=\mathcal{H}\big(f_{\bm{\theta}\odot\bm{m}_{\alpha}}(\bm{x})\big)$
+19:   Select task index via gradient signal:
+20:   $t_{\bm{x}}=\arg\min_{t}\frac{\partial\mathcal{H}}{\partial\alpha_{t}}$
+21:  Prompt selection
+22:  Select E-Prompt $\bm{e}_{t_{\bm{x}}}$ and soft-network $\bm{m}_{t_{\bm{x}}}$
+23:  Construct prompted model $f_{\bm{g},\bm{e}_{t_{\bm{x}}},\bm{m}_{t_{\bm{x}}}}$:
+24:   Attach $\bm{g}$ at layers $l^{start}_{g},l^{end}_{g}]$
+25:   Attach $\bm{e}_{t_{\bm{x}}}$ at layers $l^{start}_{e},l^{end}_{e}]$
+26:   Apply soft-transformer routing via $f_{\bm{\theta}\odot\bm{m}_{t_{\bm{x}}}}$
+27:  Prediction:
+28:   $\hat{y}=f_{\bm{g},\bm{e}_{t_{\bm{x}}},\bm{m}_{t_{\bm{x}}}}(\bm{x})$
+Table 7: Comparisons with others on 10-Split-CIFAR100.
+10-Split-CIFAR100
+Method ACC($\uparrow$) Forget($\downarrow$)
+L2P 83.38 5.88
+DualPrompt 86.61 5.86
+ESN 86.42 6.08
+CODA-Prompt 85.73 7.13
+CPrompt 77.14 5.97
+DualPrompt+Soft-TF 92.35 2.98
+Additional Results on CUB-200 & 5-Datasets.
+As shown in Table˜8, combining NoRGa with Soft-TF yields consistently superior fine-tuning performance across both Split CUB-200 and 5-Datasets under different pretraining regimes.
+While NoRGa already demonstrates strong continual adaptation by mitigating task interference, the integration of Soft-TF further unlocks the expressive capacity of the frozen Transformer backbone by directly modulating attention projections through task-specific soft subnetworks.
+This results in a substantial and consistent accuracy gain over NoRGa alone, achieving near-ceiling performance on all benchmarks (e.g., 97.90% on Split CUB-200 and 95.68% on 5-Datasets with Sup-21K pretraining).
+The improvements are not marginal but systematic, indicating that Soft-TF provides a powerful complementary mechanism to NoRGa by enabling fine-grained, task-aligned parameter adaptation beyond prompt- or gradient-level regularization.
+These results clearly demonstrate that NoRGa+Soft-TF delivers an overwhelmingly strong fine-tuning strategy, effectively bridging robust continual learning with highly expressive task-specific specialization.
+Table 8: Final average accuracy (FA) on Split CUB-200 and 5-Datasets.
+Method Split CUB-200 5-Datasets
+Sup-21K iBOT-21K Sup-21K iBOT-21K
+L2P 75.46 46.60 81.84 82.25
+DualPrompt 77.56 45.93 77.91 68.03
+S-Prompt 77.13 44.22 86.06 77.20
+CODA-Prompt 74.34 47.79 64.18 51.65
+HiDe-Prompt 86.56 78.23 93.83 94.88
+NoRGa 90.90 80.69 94.16 94.92
+NoRGa + Soft-TF 97.90 90.82 95.68 95.42